### v2

# Asymptotic Bayes Optimality for Sparse Count Data

Abstract
Consider a situation of analyzing high-dimensional count data containing an excess of near-zero counts with a small number of moderate or large counts. Under the assumption that the observations are modeled by a Poisson distribution, we are interested in simultaneous testing of whether the mean of the $i^{\text{th}}$ observation is small or large. In this work, we study some optimal properties (in terms of Bayes risk) of multiple-testing rules when the mean parameter is modeled by both two-group and a general class of one-group shrinkage priors. The class of one-group priors was proposed by [Polson and Scott (2010)] and studied by [Ghosh et al. (2016)] in the context of the normal means model and by [Tang et al. (2018)] for the linear regression model. Here, first, we model each mean by a two-group prior, and under additive 010-1 loss function, obtain an expression for the optimal Bayes risk under some assumption similar in the spirit of [Bogdan et al. (2011)]. Next, assuming that the observations are truly generated from a two-group mixture model and modelling each mean parameter by the broad class of one-group priors, we study the Bayes risk induced by our chosen class of priors. We have been able to show that, when the underlying level of sparsity is known, under some proposed assumptions, the Bayes risk corresponding to our broad class of priors attains the optimal Bayes risk, upto a multiplicative constant. When this sparsity pattern is unknown, motivated by [Yano et al. (2021)], we use an empirical Bayes estimate of the global shrinkage parameter. In this case, also, we show that the modified decision rule attains the optimal Bayes risk, upto a multiplicative constant. In this way, as an alternative solution for two-group prior, we propose a broad class of global-local priors having similar optimal properties in terms of Bayes risk for quasi-sparse count data. Finally, the theoretical results are verified using simulation studies followed by a real data analysis.
1 Introduction
In this era of enormous supply of data with a large number of variables, multiple hypothesis testing has become an area of keen interest for researchers in statistics. One of the key features in many high-dimensional situations, i.e., where the number of variables/parameters grow at least as fast as the sample size, is that the number of significant variables/parameters is rather small compared to the number of all possible variables/parameters. This is the so-called phenomenon of sparsity at a broad level and such data appear in the context of various scientific investigations, see, e.g., [Buhlmann, and Van De Geer (2011)], [Mitchell and Beauchamp (1988)], and [Johnstone and Silverman (2004)]. The theoretical investigation of multiple testing rules in sparse high-dimensional settings has been a major focus of research over the last decade and a half, see, e.g., [Bogdan et al. (2008)], [Bogdan et al. (2011)], [Datta and Ghosh (2013)], [Carvalho et al. (2009)], [Carvalho et al. (2010)], [Neuvial and Roquain (2012)] and [Tang et al. (2018)].
    Most of these developments, however, have focused on data generated from normal distributions (e.g., observations coming from the normal means model or observations modelled by normal linear regression) or some related continuous distributions. However, the assumption of normality is not reasonable for many types of data and for that matter, statisticians often face problems where assumption of continuity of distribution becomes untenable. There is a need, therefore, to develop and theoretically investigate multiple testing methods in such problems and such problems can be of great practical relevance as specific examples will bear out shortly. Our work in this paper is a modest step in that direction. The existing literature in this domain is rather scarce but these limited sources inspired us to undertake this study. Specifically, in this paper, we focus on proposing and theoretically investigating multiple hypothesis testing rules for parameters of models for count (of certain events) data. The description, background and motivation of our work follow in the next few paragraphs.
    Suppose one is interested in modelling the counts of an event, for example, a terrorist attack, happening in several areas of a country or several countries in a continent over a certain period of time. A common finding in such data is that most of the counts are close to zero (but not necessarily exactly zero) while some are moderate or large. This would usually indicate that the actual “rates” of occurrences are small for most areas, while a relatively few areas may be seriously prone to such attacks. Our interest in such data is partly motivated by the works of [Datta and Dunson (2016)] and [Hamura et al. (2022)], who dealt with similar settings. [Datta and Dunson (2016)] referred to data of the kind just described as “quasi-sparse” count data and their paper studied, among other things, multiple testing in the context of such data. [Datta and Dunson (2016)] and [Hamura et al. (2022)] assumed that the $i$-th observation follows a Poisson distribution with mean $\theta_{i}$, i.e., if the observation vector is $\mathbf{Y}=(Y_{1},Y_{2},\cdots,Y_{n})$, then $Y_{i}\stackrel{{\scriptstyle ind}}{{\sim}}Poi(\theta_{i})$, where $Poi(\theta_{i})$ denotes a Poisson distribution with mean $\theta_{i}$. A question of interest (e.g., in the context of data on terrorist attacks or any crime for that matter) would be to identify the ‘hotspots’ i.e. areas with large $\theta_{i}$’s. This can be framed as a multiple testing problem for deciding which $\theta_{i}$’s are actually large and which are small. [Datta and Dunson (2016)] and [Hamura et al. (2022)] followed a Bayesian viewpoint and modelled $\theta_{i}$’s by appropriate priors. As mentioned in [Datta and Dunson (2016)], efforts for estimating the $\theta_{i}$’s corresponding to Poisson counts data back to the empirical Bayes approaches of [Robbins (1956)] and [Kiefer and Wolfowitz (1956)]. It may be noted that [Datta and Dunson (2016)] is specifically focused on efficient handling of quasi-sparse count data, contrary to the rather general appeal of [Robbins (1956)] and [Kiefer and Wolfowitz (1956)]. Since in technical and conceptual aspects, the present paper is mainly motivated by [Datta and Dunson (2016)], we will not delve into further discussions on the above two references. We first present a summary of the approach of [Datta and Dunson (2016)]. This will be followed by a description of our work.
    The most important challenge of modelling quasi-sparse count data is to be able to take into account the great abundance of near zero counts and the simultaneous occurrences of some moderate and large counts. A common approach of modelling quasi-sparsity is to employ standard models dealing with “zero-inflation”. In the Bayesian setting, this would amount to using some form of mixture prior for the mean parameters. As mentioned by [Datta and Dunson (2016)], one possible such choice is mixture of a distribution degenerate at zero with either another distribution degenerate at some non-zero value or a Gamma distribution, the mixture components arising with probabilities 1p1-p and $p$, respectively, for some small $p$. As a result, the marginal distribution of $Y_{i}$ becomes a mixture of a degenerate distribution at zero with a Poisson or a negative binomial distribution. See [Yang et al. (2009)] in this context. As commented by [Datta and Dunson (2016)], such mixture modelling may be computationally unstable and also unsatisfactory in some sense in terms of performance, as illustrated in their simulations. [Datta and Dunson (2016)] instead proposed to model data of such kind through a unimodal continuous prior on $\theta_{i}$. Specifically, they considered the following hierarchical prior:
$\theta_{i}|\lambda_{i},\tau\stackrel{{\scriptstyle ind}}{{\sim}}{}Ga(\alpha,\lambda^{2}_{i}\tau^{2}),\lambda^{2}_{i}\stackrel{{\scriptstyle ind}}{{\sim}}{}\pi_{1}(\lambda^{2}_{i}),\tau^{2}\sim\pi_{2}(\tau^{2}),$ (1)
where $Ga(\alpha_{1},\alpha_{2})$ denotes a Gamma distribution with shape parameter $\alpha_{1}(>0)$ and scale parameter $\alpha_{2}(>0)$ with p.d.f. $f(x|\alpha_{1},\alpha_{2})=\frac{1}{\Gamma(\alpha_{1}){\alpha}^{\alpha_{1}}_{2}}x^{\alpha_{1}-1}e^{-x/\alpha_{2}}$ for $x>0,$ while $\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$ are densities for $\lambda^{2}_{i}$ and $\tau^{2}$, respectively. Here $\lambda_{i}$’s are known as the local shrinkage parameters and $\tau$ is termed the global shrinkage parameter. This is an example of application of an one-group global-local shrinkage prior for modelling the means of independent Poisson distributions. In this hierarchical formulation, $\tau$ is used to induce an overall (hence global) sparsity and therefore a small value of it ensures a large mass of the distribution of $\theta_{i}$ around the origin. On the other hand, $\lambda_{i}$’s are intended to act “locally” to modulate the shrinkage corresponding to individual parameters and hence the prior on them should have thick tails in order to ensure that the observations corresponding to true signals are left mostly unshrunk but the noises are indeed shrunk towards zero. See [Carvalho et al. (2009)] and [Carvalho et al. (2010)] in this context. It may be recalled that in the context of normal data, continuous one-group shrinkage priors have been used extensively, see e.g., [Carvalho et al. (2009)], [Carvalho et al. (2010)], [Datta and Ghosh (2013)], [Ghosh et al. (2016)], [Paul and Chakrabarti (2025)], just to name a few. Some examples of one-group global-local shrinkage priors are the three parameter beta normal prior of [Armagan et al. (2011)] (which contains the horseshoe of [Carvalho et al. (2009)], [Carvalho et al. (2010)]), the generalized double Pareto prior of [Armagan et al. (2013)], and the normal-exponential-gamma prior of [Brown and Griffin (2010)]. They were originally proposed as alternatives to their two-group counterparts in sparse problems under normality since such priors can substantially reduce computational complexity of doing inference using two-group priors. See, in this context, the discussions in [Carvalho et al. (2010)]. As far as we know, [Datta and Dunson (2016)] were the first to extend the application of one-group global-local shrinkage priors in the context of non-normal data with sparse parameter vectors.
    For proving their theoretical results, [Datta and Dunson (2016)] either used $\tau$ in (1) as fixed or let $\tau\to 0$ as $n\to\infty$. Integrating out $\theta_{i}$ and defining $\kappa_{i}=1/(1+\lambda^{2}_{i}\tau^{2})$, the marginal distribution of $Y_{i}$ given $\kappa_{i}$ is obtained as
$f(Y_{i}|\kappa_{i})\propto(1-\kappa_{i})^{Y_{i}}\kappa^{\alpha}_{i},$
i.e., the marginal distribution of each $Y_{i}$ (given $\kappa_{i}$) follows a negative binomial distribution with size $\alpha$ and probability of success $1-\kappa_{i}$. We note that $\theta_{i}|Y_{i},\kappa_{i}\sim Ga(Y_{i}+\alpha,1-\kappa_{i}),$ and $\mathbb{E}(\theta_{i}|Y_{i},\kappa_{i})=(1-\kappa_{i})(Y_{i}+\alpha)$. Then $(1-\kappa_{i})$ is the factor by which $(Y_{i}+\alpha)$ is shrunk towards zero in the formula of the posterior mean. This is similar to shrinkage of usual MLE towards zero in the expressions for Bayes estimates in normal models. [Datta and Dunson (2016)] assumed the Gauss Hypergeometric (GH) prior on $\kappa_{i}$ given by:
$\pi(\kappa_{i}|a_{1},a_{2},\tau,\gamma)=C_{2}{\kappa}^{a_{1}-1}_{i}(1-\kappa_{i})^{a_{2}-1}[1-(1-\tau^{2})\kappa_{i}]^{-\gamma},0<\kappa_{i}<1,a_{1},a_{2},\tau,\gamma>0\hskip 1.42271pt,$ (2)
where $C^{-1}_{2}=Beta(a_{1},a_{2})_{2}F_{1}(\gamma,a_{1},a_{1}+a_{2},1-\tau^{2})$ is the normalizing constant, $Beta(a_{1},a_{2})=\int_{0}^{1}t^{a_{1}-1}(1-t)^{a_{2}-1}dt$ is the usual beta function and $_{2}F_{1}$ the Gauss hypergeometric function (see [Armero and Bayarri (1994)] for the definition). For their theoretical analysis, [Datta and Dunson (2016)] further fixed $a_{1}=a_{2}=\frac{1}{2}$. The choice of the GH prior is motivated by the additional flexibility in the prior mass (of $\kappa_{i}$) brought in the regions near 0 and $1$ leading to flexible shrinkage; see [Datta and Dunson (2016)] for more detail. Part of this can be explained by the flexibility in choosing $\gamma$. They established that for their proposed hierarchical form (1), with $\kappa_{i}$ modelled as (2), the marginal prior density of $\theta_{i}$ given $\tau$ is unbounded at origin, i.e., $\theta_{i}$ has a pole at zero. This ensures high probability near zero in the marginal distribution of $Y_{i}$. The abundance of near-zero counts in the data observed is thus well explained by the prior. They also showed that, for their proposed class of priors, the posterior distribution of $\kappa_{i}$ concentrates near 0 or $1$ depending on whether the observation is large or small, thus ensuring minimal shrinkage or heavy shrinkage for large or small observations, respectively.
    As indicated before, in the Bayesian paradigm, the usual approach to model a vector $\boldsymbol{\theta}$ with mostly small values of coordinates is through a two-group prior like a spike-and-slab prior by mixing a distribution concentrated around 0 with a heavy-tailed distribution (for capturing the large values), the mixture components appearing with probabilities (1p)(1-p) and $p$ respectively with a small $p$. [Datta and Dunson (2016)] remarked that one natural choice of a two-group mixture prior to incorporate quasi-sparsity for Poisson counts is given by:
$\theta_{i}\stackrel{{\scriptstyle iid}}{{\sim}}(1-p)Ga(\alpha,\beta)+pGa(\alpha,\beta+\delta),i=1,2,\cdots,n.$ (3)
By keeping $\alpha$ and $\beta$ small one can ensure high concentration of a $Ga(\alpha,\beta)$ distribution near zero while letting $\delta>>\beta$ ensures that $Ga(\alpha,\beta+\delta)$ is flatter than the $Ga(\alpha,\beta)$ distribution. For such choices of $\alpha,\beta$ and $\delta$, $Ga(\alpha,\beta)$ is a natural prior for modelling the small $\theta_{i}$’s while $Ga(\alpha,\beta+\delta)$ is a natural prior for modelling the large or moderate $\theta_{i}$’s. Note that, using (3), the marginal distribution of $Y_{i}$ is obtained as a mixture of two negative binomial distributions having the form:
$Y_{i}\stackrel{{\scriptstyle iid}}{{\sim}}(1-p)NB(\alpha,\frac{1}{\beta+1})+pNB(\alpha,\frac{1}{\beta+\delta+1}),i=1,2,\cdots,n.$ (4)
Here $NB(\alpha,p)$ denotes a negative binomial distribution with size $\alpha$ and probability of success $p$. It may be noted that, by introducing a set of latent variables $\nu_{i},i=1,2,\cdots,n$, such that $\nu_{i}=0$ indicates $H_{0i}:\theta_{i}\sim Ga(\alpha,\beta)$ is true and $\nu_{i}=1$ indicates $H_{1i}:\theta_{i}\sim Ga(\alpha,\beta+\delta)$ is true with $P(\nu_{i}=1)=p$ and $P(\nu_{i}=0)=1-p$, the marginal prior on $\theta_{i}$ becomes of the form (3). [Datta and Dunson (2016)] proposed an interesting multiple testing rule (described shortly) for testing $H_{0i}$ vs $H_{1i}$ for $i=1,2,\cdots,n$ simultaneously, using their one-group priors of the form (1) and (2) when the true $\theta_{i}$’s are generated from (3) for Poisson count data. Before going into the details we briefly describe in the next paragraph the motivation behind this rule.
    [Carvalho et al. (2010)] in their seminal paper on horseshoe prior (in the context of normal means model) observed a very interesting fact. They compared the expression for the posterior mean under a spike-and-slab two-group model (with a heavy tailed slab part) with that under the horseshoe. This comparison indicates that in sparse situations, the expected posterior shrinkage coefficient $\mathbb{E}(1-\kappa_{i}|\text{data})$ plays approximately the same role as the posterior inclusion probability $\mathbb{P}(\theta_{i}\neq 0|\text{data})$ corresponding to a two-group model. It is standard fact that when $\theta_{i}$’s are generated from a two-group model, the optimal multiple testing rule for testing $H_{0i}$ vs $H_{1i}$ for $i=1,2,\cdots,n$ simultaneously, with respect to an additive 010-1 loss function, would reject $H_{0i}$ in favor of $H_{1i}$ if $\mathbb{P}(\theta_{i}\neq 0|\text{data})>\frac{1}{2}$. These two facts inspired [Carvalho et al. (2010)] to propose a multiple testing rule based on horseshoe prior that rejects $H_{0i}$ if $\mathbb{E}(1-\kappa_{i}|\mathbf{X})>\frac{1}{2}$ and apply the rule when $\theta_{i}$’s are truly generated from a two-group mixture prior. [Carvalho et al. (2010)] then conducted a simulation study assuming that observations ($\mathbf{X}=(X_{1},\cdots,X_{n})$) are coming from a sparse normal means model ($X_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(\theta_{i},1),i=1,2,\cdots,n$) with different sparsity levels while the true signals are fixed at equispaced points ranging from $0.5$ to $5.0$. They then applied both the horseshoe prior ($\theta_{i}|\lambda_{i},\tau\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(0,\lambda^{2}_{i}\tau^{2}),\lambda_{i}\stackrel{{\scriptstyle iid}}{{\sim}}C^{+}(0,1),\tau\sim C^{+}(0,1)$) and the two-group prior for the parameters $\theta_{i}$’s and obtained simulation estimates of certain posterior measures corresponding to these two types of priors, based on the same data sets. The estimated shrinkage weights resulting from the horseshoe prior (i.e., the posterior mean of $(1-\kappa_{i})$ given $\mathbf{X}$ and $\tau$) were very close to the estimated inclusion probabilities corresponding to their two-group counterparts, i.e., $\mathbb{P}(\theta_{i}\neq 0|\mathbf{X})$, for a wide range of sparsity patterns. So the simulations corroborated their intuition that $\mathbb{E}(1-\kappa_{i}|\text{data})$ and $\mathbb{P}(\theta_{i}\neq 0|\mathbf{X})$ play the same role approximately in sparse situations. They also observed that the simulation performances of the optimal rule for the two-group setting and their proposed rule based on horseshoe prior were close in terms of misclassification probability. It was then natural for researchers to ask whether the good simulation results indeed had any formal decision theoretic explanation. [Datta and Ghosh (2013)] studied the problem and proved that under additive 010-1 loss function, the decision rule based on horseshoe prior attains the optimal Bayes risk, upto a multiplicative constant, when the data is truly generated from a two-group model. Motivated by this result, for the normal means model, [Ghosh et al. (2016)] considered a more general class of priors (originally suggested by [Polson and Scott (2010)]) where the local shrinkage parameters are modelled as:
$\pi_{1}(\lambda^{2}_{i})=K(\lambda^{2}_{i})^{-a-1}L(\lambda^{2}_{i}),$ (5)
where $K>0$ is the constant of proportionality and $a>0$ and $L:(0,\infty)\to(0,\infty)$ is a measurable non-constant slowly varying function in Karamata’s sense (see [Bingham et al. (1987)]), that is, $\frac{L(\alpha x)}{L(x)}\to 1$ as $x\to\infty$, for any $\alpha>0$. They were able to establish that the similar optimality results hold, in fact with better constants, for their choice of broad class of global-local priors containing horseshoe as a special case.
    The above positive results about the optimal performance of one-group priors in two-group settings inspired [Datta and Dunson (2016)] to investigate the performance of a multiple testing rule for the quasi-sparse count data (for testing $H_{0i}$ vs $H_{1i}$ simultaneously for $i=1,2,\cdots,n$) based on GH prior (due to its computational amenability compared to its two-group counterpart stated earlier). Their rule rejects $H_{0i}$ if $1-\mathbb{E}(\kappa_{i}|Y_{i},\alpha,\tau,\gamma)>\zeta$, where $\zeta$ is a suitably chosen threshold. The main intuition behind this comes from a simple comparison of the expressions of the posterior means under the two-group model (3) and their one-group prior as done previously in [Carvalho et al. (2010)] in the context of normal models. This is explained in Section 3. In a simulation study, [Datta and Dunson (2016)] showed that the performance of their proposed decision rule using GH prior with $a_{1}=a_{2}=\frac{1}{2},\gamma\neq 1$ is better than that of the horseshoe prior ($a_{1}=a_{2}=\frac{1}{2},\gamma=1$) for simultaneous testing of $H_{0i}$ vs $H_{1i}$ for $i=1,2,\cdots,n$, in terms of the misclassification probability when the data is generated from a two-group mixture. They also showed that simulation estimate of the average mean squared error of the estimate of $\theta_{i}$’s based on the GH prior with $\gamma\neq 1$ is smaller than that using the horseshoe estimator $(\gamma=1)$. They also stated an upper bound on the type I error of their decision rule when $\theta_{i}$’s are truly generated using (3). However, they did not obtain any expression for the type II error or that of the Bayes risk based on GH prior. They also did not study the optimal Bayes risk in the two-group formulation (3) based on any decision theoretic framework and for that matter, this study has not been undertaken in the literature till now by any other authors, in the context of count data, as far as we know.
    The above discussion brings forth many interesting questions which warrant further study. Our work is a modest attempt to address some of these. First and foremost, it would be of interest to find expressions for the optimal Bayes risk with respect to some natural loss functions for multiple hypothesis testing for the $\theta_{i}$’s when they are indeed generated from a two-group gamma mixture as in (3). Secondly, it would be of keen interest to investigate if other choices of one-group shrinkage priors can be employed for simultaneous testing on quasi-sparse count data and if they can match (or better) the good performance/properties of inference using GH prior, as shown in this context in [Datta and Dunson (2016)]. Most importantly, it will be very interesting to settle, if one-group priors can indeed be proxies of their two-group counterparts when evaluated in a decision theoretic sense and if they can attain the corresponding optimal Bayes risk, even in the context of count data of this kind. Inspired by the results of [Carvalho et al. (2009)], [Carvalho et al. (2010)], [Ghosh et al. (2016)] and [Datta and Dunson (2016)], we have considered in this paper a broad class of global-local shrinkage priors of the form (1) with the local shrinkage parameter $\lambda_{i}$ modelled as (5), where $a\geq 1$, and $K$ and $L(\cdot)$ are same as before. For $a\geq 1$ in (5), this class of priors includes big subclasses of three parameter beta normal priors introduced by [Armagan et al. (2011)] and the generalized double Pareto prior due to [Armagan et al. (2013)]. These priors are discussed in detail in Section 3.
   We now briefly highlight the main contents of this paper and our contributions. We start with our theoretical results. In Section 2, under the additive symmetric 010-1 loss function, we have obtained an asymptotic expression for the Bayes risk (in Theorem 1) corresponding to the optimal rule (also called the Bayes oracle rule) induced by the two-group prior given in (3) under some assumptions on the model parameters $p,\beta$ and $\delta$. This asymptotic framework is inspired by [Bogdan et al. (2011)] and [Ghosh et al. (2016)]. The conditions on the parameters in this framework are somewhat similar in spirit to those introduced by [Bogdan et al. (2011)]. However, since both the level of sparsity $p$ and the scale parameter $\beta$ under the null go to zero, an extra assumption on the relative growth of $\frac{\beta}{p}$ becomes necessary. These conditions are discussed in detail in Section 2. In Section 3, as an alternative to the two-group approach, we propose a multiple testing rule based on our chosen class of one-group priors to be used for simultaneous testing when the true $\theta_{i}$’s are generated from a two-group model. The rule is similar to that in [Datta and Dunson (2016)] and is defined in (16). Corresponding to our class of priors, we provide some posterior concentration inequalities for $\kappa_{i}$ in Subsection 4.2 (see Theorems 3-6). The conclusions about concentration and shrinkage properties obtained in these results are similar to those obtained by [Ghosh et al. (2016)] for the sparse normal means model and [Datta and Dunson (2016)] for quasi-sparse count data. Based on the concentration inequalities, in Subsection 4.3, we obtain non-trivial upper bounds for both type I and type II error probabilities of this rule. These are reported in Theorems 10 and 12 respectively. When the theoretical proportion of non-nulls ($p$) is known, by suitably fixing the global shrinkage parameter $\tau$ in terms of $p$, we prove that the Bayes risk corresponding to the decision rule (16) based on our proposed class of priors attains the optimal Bayes risk, upto some multiplicative constant. This is the first main contribution of this work and is stated in Theorem 1. When the theoretical proportion of non-nulls is unknown, based on an empirical Bayes estimate $\widehat{\tau}$ of $\tau$ as proposed by [Yano et al. (2021)], in Subsection 4.4, we provide bounds on the probabilities of type I and type II errors of our rule where $\tau$ is replaced by $\widehat{\tau}$. These are presented in Theorems 14 and 15 respectively. We are able to show that the Bayes risk of the this empirical Bayes decision rule also attains the optimal Bayes risk, upto some multiplicative constant. This is formally stated in Theorem 2. These results are the first of their kind in the literature.
    The conclusions of our main results are similar to those of Theorem 1 and Theorem 2 of [Ghosh et al. (2016)]. It must be stressed that these results are not obtained by merely following the techniques of [Ghosh et al. (2016)]. Non-trivial modifications of different arguments used by them are necessary. It may be noted that Theorems 1 and 2 of our paper are based on a series of lemmas about several quantities related to the posterior distribution of the shrinkage coefficient, leading to bounds on Type I and Type II errors. While proving Theorem 12 (bounding Type II error), we have used some arguments similar to those of Theorem 6 of [Ghosh et al. (2016)]. However, Theorem 10 (related to Type I error) derived here is completely independent of the work of [Ghosh et al. (2016)]. This proof hinges upon the use of Theorem 4. See Remark 5 in Section 4.2 in this context. Similarly, Theorems 14 and 15 (dealing with the empirical Bayesian rule) require invoking arguments substantially different from these employed for proving similar results in [Ghosh et al. (2016)]. See in this context Remark 16 of Section 4.4.
    As far as we know, this is the first formal study of decision theoretic optimality of one-group priors in the context of sparse count data. Our paper hopefully presents a strong case for application of one-group priors in this context as an alternative to two-group mixture priors. We should mention that since GH prior is not included in our chosen class of priors, our results do not provide any theoretical guarantee that the decision rule based on GH prior also attains the optimal Bayes risk, upto some multiplicative constant. However, in our simulation studies, we have observed that performance of testing rules based on our prior and GH prior are quite similar to that of the optimal rule in two-group setting in terms of misclassification probability for a wide range of sparsity levels, specifically for small values of $p$. Our method has also been applied for the analysis of a real data set and has returned pretty satisfactory results. The simulation results and the real-data application are presented in Sections 5 and 6, respectively. Our simulation (see Table 1) also highlights that the role of the actual choice of $\tau$ for a given $p$ is very important to ensure good match between the asymptotic risk of the Bayes Oracle and our method. In Section 7, we present an overall discussion along with some possible extensions of this work. Section 8 contains the proofs of all theorems and other theoretical results.
1.1 Notation
For any two sequences of real numbers $\{a_{n}\}$ and $\{b_{n}\}$ with $b_{n}\neq 0$ for all $n$, $a_{n}\sim b_{n}$ implies $\lim_{n\to\infty}a_{n}/b_{n}=1$. By $a_{n}=O(b_{n})$, and $a_{n}=o(b_{n})$ we denote $|a_{n}/b_{n}|<M$ for all sufficiently large $n$, and $\lim_{n\to\infty}a_{n}/b_{n}=0$, respectively, $M>0$ being a global constant that is independent of $n$. Likewise, for any two positive real-valued functions $f_{1}(\cdot)$ and $f_{2}(\cdot)$ with a common domain of definition that is unbounded to the right $f_{1}(x)\sim f_{2}(x)$ denotes $\lim_{x\to\infty}f_{1}(x)/f_{2}(x)=1.$ Throughout this article, the indicator function of any set $A$ will always be denoted $\mathbf{1}\{A\}$.
2 The Two-group prior and the Bayes Oracle
To capture the sparsity of the $\boldsymbol{\theta}$ in the context of our problem, the popular approach would be to use a two-component mixture prior where with probability $p$, $\theta_{i}$ is degenerate at zero and with probability (1p)(1-p) it is generated either from a distribution degenerate at some non-zero value or a Gamma distribution. See [Datta and Dunson (2016)] and the references therein in this context. But as commented by them, such mixture modelling may cause computational instability. Inspired by the work of [Datta and Dunson (2016)], we model each $\theta_{i}$ as a sample from a scale mixture of two gamma distributions of the form (3). As a result of this, the marginal distribution of the observations is a mixture of two negative binomial distributions as given in (4).
    Now we consider the problem of multiple hypothesis testing of $H_{0i}:\nu_{i}=0$ against $H_{1i}:\nu_{i}=1$, for $i=1,2,\cdots,n$ within a Bayesian decision theoretic framework. We consider the usual 010-1 loss function for individual tests and assume that the overall loss of a multiple testing procedure is the sum of losses corresponding to individual tests. In this way, our approach is based an additive loss function and is similar to that of [Bogdan et al. (2011)] for the normal means model. Additive losses were earlier considered by [Lehmann (1957)] in the context of testing.
    Let $t_{1i}$ and $t_{2i}$ denote the probabilities of type I and type II error of the $i^{\text{th}}$ test respectively, and are defined as
$t_{1i}=P_{H_{0i}}(H_{0i}\hskip 1.42271pt\text{is}\hskip 1.42271pt\text{rejected})$
and
$t_{2i}=P_{H_{1i}}(H_{0i}\hskip 1.42271pt\text{is}\hskip 1.42271pt\text{accepted}).$
Then the Bayes risk $R$ of a multiple testing procedure is obtained as
$R=\sum_{i=1}^{n}[(1-p)t_{1i}+pt_{2i}].$ (6)
Recalling [Bogdan et al. (2011)], it is easy to see that a multiple testing rule minimizing the Bayes risk $R$ is the one that applies the simple Bayes classifier for each individual test. So the optimal rule, for each $i$, rejects $H_{0i}$ in favor of $H_{1i}$ if
$\frac{f_{H_{1i}}(Y_{i})}{f_{H_{0i}}(Y_{i})}>\frac{1-p}{p}\hskip 1.42271pt,$ (7)
where $f_{H_{1i}}$ and $f_{H_{0i}}$ are the marginal densities of $Y_{i}$ under the alternative and null hypotheses respectively. Using (4), the optimal testing rule given in (7) is simplified as
$\text{reject}\hskip 1.42271ptH_{0i}\hskip 1.42271pt\text{if}\hskip 14.22636ptY_{i}>C\hskip 1.42271pt,$ (8)
where
$C=C_{p,\alpha,\beta,\delta}=\frac{\log(\frac{1-p}{p})+\alpha\log(\frac{\beta+\delta+1}{\beta+1})}{\log[\frac{(\beta+\delta)(\beta+1)}{\beta(\beta+\delta+1)}\bigg{]}}\hskip 1.42271pt.$ (9)
Due to the presence of $p,\alpha,\beta$ and $\delta$, the decision rule (8) is termed as Bayes Oracle. Its performance can not be equalled by any testing rule based on finite samples if these parameters are unknown. Let us denote the Bayes risk of the Bayes Oracle as $R^{\text{BO}}_{\text{Opt}}$.
    Note that, under the mixture model (4), the marginal distributions of $Y_{i}$ under both null and alternative hypotheses are independent of the choice of $i$. The threshold of the Bayes Oracle, given in (9), corresponding to a two-group prior is also independent of $i$. Therefore, we drop $i$ from $t_{1i}$ and $t_{2i}$ and rename those as $t^{\text{BO}}_{1}$ and $t^{\text{BO}}_{2}$, respectively. Therefore, we can write
$R^{\text{BO}}_{\text{Opt}}=n[(1-p)t^{\text{BO}}_{1}+pt^{\text{BO}}_{2}].$ (10)
As in [Bogdan et al. (2011)], we want to study the behaviour of $R^{\text{BO}}_{\text{Opt}}$ under a suitable asymptotic setting. Since our basic model is non-normal, we cannot simply mimic the setting of [Bogdan et al. (2011)] as in their Assumption (A). Before writing out the asymptotic setting, we want to motivate the logic behind it. Since we are dealing with a sparse situation and a two-group Gamma mixture prior aimed at capturing quasi-sparse counts, we first study the behaviour of $t_{1}$ and $t_{2}$ under the basic assumptions $p\to 0$, $\beta\to 0$, and $\delta\to\infty$ as $n\to\infty$. Calculations of Theorem 1 reveal that under these assumptions, the error bounds corresponding to the Bayes oracle rule (8) (with $C$ as in (9)) are of the form,
$t^{\text{BO}}_{1}$ $\leq\alpha\cdot\frac{\beta\log(\frac{1}{\beta})}{\log(\frac{1}{p})}(1+o(1))$
and
$t^{\text{BO}}_{2}$ $=P_{H_{1i}}(Z_{i}\leq\frac{\log(\frac{1}{p})}{\sqrt{\alpha}\delta\log(\frac{1}{\beta})}(1+o(1))-\sqrt{\alpha}(1+o(1)))\hskip 1.42271pt,$
where $Z_{i}$ is the standardized version of $Y_{i}$ under $H_{1i}$} (obtained by subtracting its mean and dividing by its standard deviation under $H_{1i}$}). We can show using the expression of $t^{\text{BO}}_{2}$ that the power of the Bayes Oracle rule is asymptotically strictly between 0 and $1$ if $\beta\propto p^{C_{1}}$ for some $C_{1}>1$, $\frac{\log\delta}{\log(\frac{1}{p})}\to\tilde{C}$ with $\tilde{C}>\alpha C_{1}$ where $\alpha(>1)$ is bounded. In this paper, we are interested in these situations, since otherwise, even the Bayes Oracle rule cannot guarantee non-trivial inference. We now state our asymptotic setting in Assumption 1 below. Our asymptotic analysis of the Bayes Oracle (8) and our proposed rule are done under Assumption 1. This is discussed in more details later in this paper.
Assumption 1. $p\to 0$, $\alpha(>1)\textrm{ is bounded}$, $\beta\propto p^{C_{1}}$ for some $C_{1}>1$ and $\delta\to\infty$ such that $\frac{\log\delta}{\log(\frac{1}{p})}\to\tilde{C}$ with $\tilde{C}>\alpha C_{1}$ as $n\to\infty$.
    Under Assumption 1, we have,
$t^{\text{BO}}_{2}$ $=P_{H_{1i}}(Z_{i}\leq-\sqrt{\alpha}(1+o_{1}(1)))$
and $t^{\text{BO}}_{1}=o(p)$ as $n\to\infty$. Combing these two, we have
$R^{\text{BO}}_{\text{Opt}}$ $=npP_{H_{1i}}(Z_{i}\leq-\sqrt{\alpha}(1+o_{1}(1))),$ (11)
where $o_{1}(1)$ is non-random, independent of index $i$ and goes to 0 as $n\to\infty$. Detailed calculations establishing the above facts on the error probabilities are available in the proof of Theorem 1 in Section 8. Our next target is to find expressions/upper bounds for $t_{1i}$ and $t_{2i}$ corresponding to our decision rules (16) and (18) based on priors (1) satisfying (5). This will help to study the optimality of our decision rules. In the next subsection, we motivate our decision rules by highlighting a connection between the posterior means of $\theta_{i}$ under two-group priors and our class of priors.
3 A multiple testing rule using one-group priors
In this section, we propose simultaneous hypothesis testing rules based on a broad class of global-local shrinkage priors to be applied in the context of quasi-sparse Poisson count data described before. We want to investigate in the following sections how they perform vis-a-vis the optimal rule, namely the Bayes Oracle when the means are actually generated from a two-group prior. We first briefly recall that [Datta and Dunson (2016)] had also applied a decision rule in this context using a subclass of the Gauss Hypergeometric (GH) priors described earlier. But they did not study its optimality properties in a decision theoretic framework. Our main interest lies in such a study using one-group priors in the context of count data. Towards that we consider a broad class of global-local shrinkage priors satisfying (1) and (5) described in Subsection 1. Incidentally, it may be noted that the a large spectrum of the Three Parameter Beta Normal (TPBN) priors (with $\alpha>0$ and $\beta\geq 1$) and Generalized Double Pareto priors (with $\alpha\geq 2$ and $\beta>0$) are contained in our general class. The forms of these two classes of priors are discussed in equations (12) and (13) of [Ghosh et al. (2016)]. The GH priors defined in contain the TPBN family. But [Datta and Dunson (2016)] only considered a subclass of that is mentioned before.
    We now introduce our decision rule based on our chosen class of priors. It may first be observed that the posterior distribution of $\theta_{i}$’s given $(Y_{1},\cdots,Y_{n},\tau)$ are independent and the posterior distribution of $\theta_{i}$ depends only on $Y_{i},\kappa_{i}$ and $\tau$. First note that using such priors, the posterior distribution of $\theta_{i}$ given $\kappa_{i},\tau$ and $Y_{i}$ is of the form
$\theta_{i}|Y_{i},\kappa_{i},\tau\sim\text{Ga}(Y_{i}+\alpha,1-\kappa_{i})\hskip 1.42271pt.$ (12)
By Fubini’s theorem, the posterior mean of $\theta_{i}$ is therefore given by
$E(\theta_{i}|Y_{i},\tau)=(1-E(\kappa_{i}|Y_{i},\tau))(Y_{i}+\alpha)\hskip 1.42271pt.$ (13)
Under two-group prior, the posterior distribution of $\theta_{i}$’s given $\textbf{Y}=(Y_{1},\cdots,Y_{n})$ are independent and the posterior distribution of $\theta_{i}$ only depends on $Y_{i}$. When $\theta_{i}$ is generated from the two-group prior (3), one has
$E(\theta_{i}|Y_{i})=[(1-w_{i})\frac{\beta}{\beta+1}+w_{i}\frac{\beta+\delta}{\beta+\delta+1}](Y_{i}+\alpha)=w^{*}_{i}(Y_{i}+\alpha)\hskip 1.42271pt,$ (14)
where $w_{i}=P(H_{1i}|Y_{i})$ denotes the posterior probability of $H_{1i}$ being true and $w^{*}_{i}$ is the observation-specific shrinkage weight. Comparing (13) and (14), it is apparent that the posterior shrinkage weight $1-E(\kappa_{i}|Y_{i},\tau)$ mimics the role of $w^{*}_{i}$ when one uses one-group prior instead of a two-group prior in a sparse situation.
    Under a two-group framework, the optimal multiple testing rule for an additive 010-1 loss, is given by:
$\text{Reject}\hskip 1.42271ptH_{0i}\hskip 1.42271pt\text{if}\hskip 14.22636ptw_{i}>\frac{1}{2}\hskip 1.42271pt,i=1,2,\cdots,n.$
When $\beta\to 0$ and $\delta\to\infty$ as $n\to\infty$, the first term in $w^{*}_{i}$ goes to zero while the second goes to $1$ and thus $w^{*}_{i}\to w_{i}$. As a result, the optimal testing rule is approximately:
$\text{Reject}\hskip 1.42271ptH_{0i}\hskip 1.42271pt\text{if}\hskip 14.22636ptw^{*}_{i}>\frac{1}{2}\hskip 1.42271pt,i=1,2,\cdots,n.$ (15)
Using this and the previous observation (connecting $1-E(\kappa_{i}|Y_{i},\tau)$ and $w^{*}_{i}$), a natural testing rule for simultaneous hypothesis testing using our class of priors can be taken as:
$\text{Reject}\hskip 1.42271ptH_{0i}\hskip 1.42271pt\text{if}\hskip 14.22636pt1-E(\kappa_{i}|Y_{i},\tau)>\frac{1}{2}\hskip 1.42271pt,i=1,2,\cdots,n.$ (16)
It may be recalled that the decision rule in [Datta and Dunson (2016)] based on GH prior is exactly similar and based on the same intuitive observation. Our decision rule is inspired by their work.
    Our theoretical results (Theorem 1) show excellent performance of (16) in terms of Bayes risk (with respect to additive 010-1 loss), if, among other things, $\tau$ is chosen proportional to $p$ if $p$ is known. When $p$ is unknown, this strategy will not work. In this case, one may expect to get good results if $\tau$ is replaced by some $\widehat{\tau}$ that mimics the proportion of non-zero means using observed data. Keeping this in mind and inspired by the estimate $\text{max}\{1,\sum_{i=1}^{n}\mathbf{1}\{Y_{i}\geq 1\}\}$ of the number of non-null effects in a Poisson count data considered by [Yano et al. (2021)], we propose to employ an empirical Bayes estimate $\widehat{\tau}$ of $\tau$ given by
$\widehat{\tau}=\text{max}\biggl{\{}\frac{1}{n},\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{Y_{i}\geq{1}\}\biggr{\}}.$ (17)
We then apply the decision rule (16) by replacing $\tau$ with $\widehat{\tau}$. Thus the new rule for our multiple testing problem is obtained as:
$\text{Reject}\hskip 1.42271ptH_{0i}\hskip 1.42271pt\text{if}\hskip 14.22636pt1-E(\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2}\hskip 1.42271pt,i=1,2,\cdots,n,$ (18)
[…]
    As mentioned before, our main interest is in the study of asymptotic optimality of rules (16) and (18) when applied in a two-group setting. Since both (16) and (18) depend on the posterior mean of $\kappa_{i}$, we first need the form of the posterior distribution of $\kappa_{i}$ given $Y_{i}$ and $\tau$. It may be noted that using (1) and (5), the posterior distribution of $\kappa_{i}$’s are independently distributed given $(Y_{1},Y_{2},\cdots,Y_{n},\tau)$. The posterior distribution of $\kappa_{i}$ depends only on $Y_{i}$ given $\tau$ and has the following form :
$\pi(\kappa_{i}|Y_{i},\tau)\propto\kappa^{a+\alpha-1}_{i}(1-\kappa_{i})^{Y_{i}-a-1}L(\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1))\hskip 2.84544pt,0<\kappa_{i}<1.$ (19)
In the next section, using (19), we provide some bounds on probabilities of type I and type II errors. They are later used to study the optimality property of the decision rule (16) for quasi-sparse count data generated from a two-group model. However, the posterior distribution of $\kappa_{i}$ given $Y_{i}$ and $\tau$ evaluated at $\tau=\widehat{\tau}$ does not have an explicit form like (19) since the empirical Bayes estimate of $\tau$ depends on the entire dataset $\mathbf{Y}=(Y_{1},Y_{2},\cdots,Y_{n})$. We need to devise new techniques for the study of optimality of (18) to take care of this issue. The technical details of this will be made clear later in this paper.
For the theoretical development of the paper, we consider slowly varying functions that satisfy Assumption 2 below.
Assumption 2.
(A1) There exists some $c_{0}(>0)$ such that $L(t)\geq c_{0}\;\forall t\geq t_{0}$, for some $t_{0}>0$, which depends on both $L$ and $c_{0}$.
(A2) There exists some $M\in(0,\infty)$ such that $\sup_{t\in(0,\infty)}L(t)\leq M$.
These assumptions are similar to those of [Ghosh and Chakrabarti (2017)].
4 Theoretical Results
This section is divided into several subsections, the first of which describes the main theoretical findings of this work. The second subsection is related to some concentration inequalities which are essential for proving our main results in this paper. The third one uses the results obtained in the former subsection to provide bounds on probabilities of type I and type II errors of decision rule (16). In the last subsection, we derive the bounds on probabilities of type I and type II errors of the empirical Bayes decision rule (18).
4.1 Results on Asymptotic Bayes Risk under sparsity using one-group priors
In this subsection, we present and discuss our main optimality results (in terms of Bayes risk with respect to additive 010-1 loss) for our proposed simultaneous testing rules when the true means are generated from a two-component mixture prior (3). Our optimality results are derived under the sparse asymptotic regime described in Assumption 1, described before. First, we consider the case where the theoretical proportion ($p$) of non-nulls is known. Using suitable choice of $\tau$ in terms of $p$ among other things, we are able to prove our first optimality result of this paper. It is presented in Theorem 1, proof of which is provided in Section 8.
Theorem 1.
Let $Y_{1},Y_{2},\cdots,Y_{n}$ be i.i.d. observations having a two-group mixture distribution of the form (4). Suppose we want to test $H_{0i}:\nu_{i}=0$ against $H_{1i}:\nu_{i}=1$, for $i=1,2,\cdots,n$ using decision rule (16) induced by the class of priors (1) satisfying (5), where $L(\cdot)$ satisfies (A1) and (A2) defined in Section 3. Also assume that $p,\beta$ and $\delta$ of the two-group model satisfy Assumption 1. Further assume that $\tau\to 0$ as $n\to\infty$ such that $\lim_{n\to\infty}\tau/p\in(0,\infty)$. Then given any constant $\tilde{K}\in(\frac{1}{\alpha},1)$ for all sufficiently large $n$, the Bayes risk (with respect to additive 010-1 loss) of the multiple testing rule (16), denoted $R_{\text{OG}}$, satisfies,
$1\leq\frac{R_{\text{OG}}}{R^{\text{BO}}_{\text{Opt}}}$ $\leq\frac{\tilde{K}\alpha}{\tilde{K}\alpha-1}.$ (20)
Theorem 1 shows that, for the broad class of priors of the form (1) satisfying (5), the ratio of the Bayes risk of the multiple testing rule (16) to that of the Bayes Oracle (in (8) and (9)) is bounded above by $\frac{\tilde{K}\alpha}{\tilde{K}\alpha-1}$, i.e.,
$R_{\text{OG}}=O(R^{\text{BO}}_{\text{Opt}}),\textrm{ as }n\to\infty.$
Thus the decision rule (16) based on the global-local priors asymptotically attains the optimal Bayes risk, upto some multiplicative constant under Assumption 1.
    In practice $p$ would be generally not known. We suspect that if in (16), $\tau$ is replaced by a $\widehat{\tau}$ that is a reasonable estimate of $p$, then we might be able to mimic the optimality result in the previous theorem. Our next theorem confirms this intuition, where we make use of the estimator (17) in replacing $\tau$ by $\widehat{\tau}$ in (16). The resulting empirical Bayesian multiple testing rule (18) is shown to possess similar optimality property in Theorem 2. We also need a very mild additional assumption that $p\propto n^{-\epsilon}$ for some $0<\epsilon<1$ which covers most cases of theoretical and practical interest. We emphasize that $\epsilon$ need not be known for our result, but we only need that $\epsilon\in(0,1)$. Proof of this result is provided in Section 8.
Theorem 2.
Let $Y_{1},Y_{2},\cdots,Y_{n}$ be i.i.d. observations having a two-group mixture distribution of the form (4). Suppose we want to test $H_{0i}:\nu_{i}=0$ against $H_{1i}:\nu_{i}=1$, for $i=1,2,\cdots,n$ using decision rule (18) induced by the class of priors (1) and (5), where $L(\cdot)$ satisfies (A1) and (A2). Also assume that $p,\beta$ and $\delta$ of the two-group model satisfy Assumption 1 with $p\propto n^{-\epsilon}$ for some $0<\epsilon<1$. Then given any constant $\tilde{K}\in(\frac{1}{\alpha},1)$, for all sufficiently large $n$, the Bayes risk (with respect to additive 010-1 loss) of the multiple testing rule (18), denoted $R^{\text{EB}}_{\text{OG}},$ satisfies,
$1\leq\frac{R^{\text{EB}}_{\text{OG}}}{R^{\text{BO}}_{\text{Opt}}}$ $\leq\frac{\tilde{K}\alpha}{\tilde{K}\alpha-1}.$ (21)
Theorems 1 and 2 are the first results of their kind in the literature of quasi-sparse count data. As far as we know, ours is the first formal study of asymptotic optimality of multiple testing in this context. [Ghosh et al. (2016)] proved a similar result for the same class of priors using an empirical Bayes estimate of $\tau$ given by [van der pas et al. (2014)] for the normal means model. Our result confirms similar phenomenon even if the marginal distribution of observations is Negative Binomial.
4.2 Posterior Concentration inequalities
In this subsection, we describe some concentration and moment inequalities for the posterior distribution of the shrinkage coefficient $\kappa_{i}$’s. These are essential for obtaining upper bounds on probabilities of type I and type II errors for decision rules (16) and (18). These upper bounds are discussed in the next subsections.
  Before stating our results here, first we motivate how these may be useful for deriving the bounds for type I and type II errors. Let $t_{1i}$ and $t_{2i}$ denote the probabilities of type I and type II errors, respectively, corresponding to the $i^{\text{th}}$ test. Recall by definition, $t_{1i}=\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\tau)>\frac{1}{2})$ and $t_{2i}=\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},\tau)\geq\frac{1}{2})$. Now, with (19) being the posterior distribution of $\kappa_{i}$ given $Y_{i}$ and $\tau$, it seems that finding exact asymptotic order of $t_{1i}$ and $t_{2i}$ is not feasible as the density involves an unknown $L$. Hence, [Datta and Ghosh (2013)] and [Ghosh et al. (2016)] (for normal means model) and [Datta and Dunson (2016)] (for count data) tried to derive some non-trivial asymptotic bounds for the same, although [Datta and Dunson (2016)] only derived this for $t_{1i}$. Following a similar strategy as in [Ghosh et al. (2016)], we will try to obtain bounds on $E(1-\kappa_{i}|Y_{i},\tau)$ and $E(\kappa_{i}|Y_{i},\tau)$ and apply them to produce bounds on $t_{1i}$ and $t_{2i}$.
Towards that, we first need the following concentration inequality on the posterior distribution of $\kappa_{i}$. Proof of this result is given in Section 8.
Theorem 3.
Suppose $Y_{i}\sim Poi(\theta_{i})$ independently for $i=1,2,\cdots,n$. Consider the one-group prior given in (1) satisfying (5), where $L(\cdot)$ satisfies (A1). Then for any fixed $\epsilon\in(0,1)$ and any fixed $\tau>0$,
$\mathbb{P}(\kappa_{i}<\epsilon|Y_{i},\tau)\leq\frac{a}{c_{0}(a+\alpha)}(K_{0}^{-a}-K_{1}^{-a})^{-1}(\tau^{2})^{a-Y_{i}}\bigg{(}\frac{\epsilon}{1-\epsilon}\bigg{)}^{a+\alpha}L\bigg{(}\frac{1}{\tau^{2}}\bigg{)}(1+K_{1}\tau^{2})^{(Y_{i}+\alpha)}(1+o(1))\hskip 1.42271pt,$
where $o(1)$ is non-random, independent of index $i$ and depends only on $\tau$ such that $\lim_{\tau\to 0}o(1)=0$, and $K_{0}$ and $K_{1}$ are as in Lemma 18.
It may be noted that this result is true by taking any $a>0$ in the definition of $\pi_{1}(\lambda^{2}_{i})$.   Next we describe an upper bound on the posterior mean of the shrinkage coefficient $(1-\kappa_{i})$. Later, this bound will be used for proving an upper bound on $t_{1i}$. Proof of the theorem is provided in Section 8.
Theorem 4.
Consider the setup of Theorem 3 with the global-local prior of the form (1) satisfying (5) where $L(\cdot)$ satisfies Assumption 2 of Section 3. Then, for any fixed $\tau>0$ and for any $Y_{i}\in[0,a-1)$,
$\mathbb{E}(1-\kappa_{i}|Y_{i},\tau)\leq\frac{a}{c_{0}}(K_{0}^{-a}-K_{1}^{-a})^{-1}\tau^{2}[K^{-1}+\frac{M}{(a-Y_{i}-1)}](1+K_{1}\tau^{2})^{(Y_{i}+\alpha)}\hskip 1.42271pt,$
where $K_{0}$ and $K_{1}$ are as in Lemma 18.
Remark 5.
For the normal means model, [Ghosh et al. (2016)] derived a result on the posterior distribution of $\kappa_{i}$ similar to Theorem 3 of this work. However, their result is based on the Dominated Convergence Theorem, whereas ours uses a lower bound on the normalizing quantity of the posterior distribution of $\kappa_{i}$ given $Y_{i}$ and $\tau$. They also obtained a result on the posterior mean of $1-\kappa_{i}$ similar to Theorem 4. Under Assumption 2 on $L(\cdot)$, they provided an upper bound on $\mathbb{E}(1-\kappa_{i}|Y_{i},\tau)$ by using the definition, $\mathbb{E}(1-\kappa_{i}|Y_{i},\tau)=\int_{0}^{1}\mathbb{P}(\kappa_{i}<\epsilon|Y_{i},\tau)d\epsilon$. The same technique can not be used in our situation as $\int_{0}^{1}(\frac{\epsilon}{1-\epsilon})^{a+\alpha}d\epsilon$ converges only when $0<a+\alpha<1$. However, for any $a\geq 1$, $a+\alpha\geq 1$ always. As a result, we need to use different argument than that of [Ghosh et al. (2016)] to obtain a non-trivial upper bound on the posterior expectation of the shrinkage coefficient $1-\kappa_{i}$ in Theorem 4.
After having a non-trivial bound on $\mathbb{E}(1-\kappa_{i}|Y_{i},\tau)$, our next step is to search for a similar bound on $\mathbb{E}(\kappa_{i}|Y_{i},\tau)$ for the purpose of handling the type II error. Since, for any $\eta\in(0,1)$, $\mathbb{E}(\kappa_{i}|Y_{i},\tau)$ can be expressed as the sum of $\mathbb{E}(\kappa_{i}\mathbf{1}\{\kappa_{i}>\eta\}|Y_{i},\tau)$ and $\mathbb{E}(\kappa_{i}\mathbf{1}\{\kappa_{i}\leq\eta\}|Y_{i},\tau)$, we need to find an upper bound for $\mathbb{P}(\kappa_{i}>\eta|Y_{i},\tau)$ to bound $\mathbb{E}(\kappa_{i}\mathbf{1}\{\kappa_{i}>\eta\}|Y_{i},\tau)$. The second term in the sum can be handled easily. The next theorem provides a concentration inequality in this context. Proof of this result is given in Section 8.
Theorem 6.
Under the set up of Theorem 3, for any fixed $\eta\in(0,1)$ and $\delta_{1}\in(0,1)$ and for all sufficiently small $\tau(>0)$,
$\mathbb{P}(\kappa_{i}>\eta|Y_{i},\tau)\leq\frac{(a+\alpha)}{Kc_{0}}{\tau}^{-2a}(\frac{1-\eta}{1-\eta\delta_{1}})^{Y_{i}}(\eta\delta_{1})^{-(a+\alpha)}\hskip 1.42271pt.$
It may be noted that this result is true by taking any $a>0$.
  Results obtained in Theorems 3-6 reveal some interesting characteristics of the posterior distribution of $\kappa_{i}$ based on our choice of priors. These results also lead to the following three corollaries, whose proofs are immediate and hence omitted.
  We first have a corollary of Theorem 3.
Corollary 7.
Under the assumptions of Theorem 3, ${P}(\kappa_{i}\geq\epsilon|Y_{i},\tau)\to 1$ as $\tau\to 0$ for any fixed $\epsilon\in(0,1)$ whenever $0\leq Y_{i}<a$.
[Datta and Dunson (2016)] also obtained a similar result for the Gauss hypergeometric prior under quasi-sparse count data. Thus for any fixed $0\leq Y_{i}<a$, for our choice of one-group global-local shrinkage priors, the posterior distribution of $\kappa_{i}$ concentrates near $1$ with high probability for small values of $\tau$.
We next have a corollary to Theorem 4.
Corollary 8.
Under the assumptions of Theorem 4, $\mathbb{E}(1-\kappa_{i}|Y_{i},\tau)\to 0$ as $\tau\to 0$ uniformly in $Y_{i}\in[0,a-1)$.
This corollary ensures that in the case of global-local priors, for small values of $\tau$, the noise observations will be shrunk towards the origin.
Theorem 6 yields the following corollary.
Corollary 9.
Under the assumptions of Theorem 6, $\mathbb{P}(\kappa_{i}>\eta|Y_{i},\tau)\to 0$ as $Y_{i}\to\infty$ for any sufficiently small $\tau>0$.
This corollary implies that for our choice of one-group global-local shrinkage priors, the posterior distribution of $\kappa_{i}$ concentrates near 0 with high probability for large values of $Y_{i}$ if $\tau$ is small enough.
4.3 Type I and Type II error bounds for testing the rule (16)
As mentioned earlier, this subsection provides some asymptotic bounds on $t_{1i}$ and $t_{2i}$ under the assumption that $p$ is known. Theorem 10 gives a non-trivial upper bound on the probability of type I error for the decision rule (16). On the other hand, Theorem 12 provides an upper bound on the probability of type II error. We may recall that [Datta and Dunson (2016)] did not study the decision theoretic optimality of their decision rule (based on GH prior) which has the same structure as (16). They only proved the probability of type I error goes to zero asymptotically. However, their argument has a soft spot and the upper bound for type I error also needs modification. We feel that proof of our result (Theorem 10) on the upper bound of type I error may be helpful in getting rid of these lacunae but we have not pursued this in this work. Proofs of these results are provided in Section 8.
Theorem 10.
Let $Y_{1},Y_{2},\cdots,Y_{n}$ be i.i.d. observations having a two-group mixture distribution of the form (4). Suppose we want to test $H_{0i}:\nu_{i}=0$ against $H_{1i}:\nu_{i}=1$, for $i=1,2,\cdots,n$ using decision rule (16) induced by the class of priors (1) satisfying (5), where $L(\cdot)$ satisfies (A1) and (A2). Also assume that $p,\beta,$ and $\delta$ of the two-group model satisfy Assumption 1. Further, assume that $\tau\to 0$ as $n\to\infty$. Then as $n\to\infty$, the probability of type I error of the decision rule (16), denoted $t_{1}$, satisfies
$t_{1i}\equiv t_{1}\leq\frac{2\alpha\beta}{a}.$
Remark 11.
Under Assumption 1, not only $t_{1}$ goes to zero as $n\to\infty$, but also, the rate of convergence of $t_{1}$ towards zero is faster than that of $p$. This fact will be used in obtaining an asymptotic expression for the Bayes risk $R_{\text{OG}}$, reported in Theorem 1.
Theorem 12.
Consider the set-up of Theorem 10. Further, assume that $\tau\to 0$ as $n\to\infty$ such that $\lim_{n\to\infty}\tau/p\in(0,\infty)$. Then given any $\eta\in(0,\frac{1}{2})$, $\delta_{1}\in(0,1)$ and $r=\frac{1-\eta}{1-\eta\delta_{1}}$, as $n\to\infty$, the probability of type II error of the decision rule (16), denoted $t_{2}$, satisfies
$t_{2i}\equiv t_{2}\leq\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq-\sqrt{\alpha}(1+o_{2}(1))\bigg{)},$
where $Z_{i}$ is the standardized version of $Y_{i}$ under $H_{1i}$ and the term $o_{2}(1)$ is non-random, independent of index $i$ and tends to zero as $n\to\infty$.
Remark 13.
Instead of assuming $\lim_{n\to\infty}\tau/p\in(0,\infty)$, even if we assume $\log\tau\sim\log p$ and keep other assumptions same, the calculations for proving Theorem 12 will go through and produce exactly the same upper bound on $t_{2}$ as obtained in Theorem 12. As a consequence of this, the expression for the Bayes risk $R_{\text{OG}}$ obtained by combining Theorem 10 and Theorem 12 will still hold as reported in Theorem 1.
4.4 Type I and Type II error bounds corresponding to an empirical Bayes procedure
In this subsection, we present asymptotic bounds on the probabilities of type I and type II errors of the individual decision rules (18) using the empirical Bayes estimate $\widehat{\tau}$ of $\tau$ given in (17). These are reported in Theorems 14 and 15 respectively. Approaches involved in the proofs of these theorems are significantly different from those employed for proving Theorem 10 and Theorem 12. Note that, when $\tau$ is used as a tuning parameter, the decision rule (16) corresponding to $i^{\text{th}}$ test depends on $i^{\text{th}}$ observation $Y_{i}$ only. On the other hand, $\widehat{\tau}$ depends on the entire dataset $(Y_{1},Y_{2},\cdots,Y_{n})$. Obviously, this necessitates serious changes in the approach. Towards that, we introduce a cut-off value that is asymptotically of the order $p$. Next, we divide the range of $\widehat{\tau}$ into two mutually disjoint parts depending on whether $\widehat{\tau}$ exceeds the cut-off or not. For one part, noting that for any fixed $y\geq 0$, $\mathbb{E}(1-\kappa|y,\tau)$ is non-decreasing in $\tau$, we can use results already reported for the case when $\tau$ is used as a tuning parameter. For the other part, the form of $\widehat{\tau}$ along with the independence of $Y_{i}$’s come into play for obtaining asymptotic bounds on the quantities of interest. This architecture of proof has been inspired by [Ghosh et al. (2016)] with appropriate modifications. We now present Theorem 14.
Theorem 14.
Let $Y_{1},Y_{2},\cdots,Y_{n}$ be i.i.d. observations having a two-group mixture distribution of the form (4). Suppose we want to test $H_{0i}:\nu_{i}=0$ against $H_{1i}:\nu_{i}=1$, for $i=1,2,\cdots,n$ using decision rule (18) induced by the class of priors (1) satisfying (5), where $L(\cdot)$ satisfies (A1) and (A2). Also assume that $p,\beta$ and $\delta$ of the two-group model satisfy Assumption 1 with $p\propto n^{-\epsilon}$, for some $0<\epsilon<1$. Then as $n\to\infty$, the probability of type I error of the decision rule (18), denoted $t^{\text{EB}}_{1i}$, satisfies
$t^{\text{EB}}_{1i}\leq\frac{2\alpha\beta}{a}+{\alpha\beta}+e^{-(2\log 2-1)(1-(\beta+\delta+1)^{-\alpha})np(1+o(1))}\hskip 1.42271pt,$
where the $o(1)$ is non-random, independent of index $i$ and tends to zero as $n\to\infty$.
Theorem 15.
Consider the set-up of Theorem 14. Then given any $\eta\in(0,\frac{1}{2})$, $\delta_{1}\in(0,1)$ and $r=\frac{1-\eta}{1-\eta\delta_{1}}$, as $n\to\infty$, the probability of type II error of the decision rule (18), denoted $t^{\text{EB}}_{2i}$ satisfies
$t^{\text{EB}}_{2i}\leq\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq-\sqrt{\alpha})(1+o_{2}(1))\bigg{)}(1+o(1)),$
where $Z_{i}$ is the standardized version of $Y_{i}$ under $H_{1i}$ and the terms $o_{2}(1)$ and $o(1)$ are non-random, independent of index $i$ and tend to zero as $n\to\infty$.
Remark 16.
While proving Theorem 14 and Theorem 15, we have used some arguments similar to that of [Ghosh et al. (2016)]. But only these arguments are not enough to establish these results. We need to use Theorem 10 and Theorem 12 too to complete the proofs in this subsection. Our work shows that with some non-trivial modifications, the techniques used by [Ghosh et al. (2016)] can be useful even if the underlying model is not normal.
5 Simulation Results
In this section, we are going to compare through simulations the performance of our decision rules (16) and (18) vis-a-vis the Bayes Oracle (8), when the data is truly generated from a two-group model and the data is modelled using a member of TPBN families as the prior on the $\theta_{i}$’s. We will also use the Gauss Hypergeometric (GH) prior of [Datta and Dunson (2016)]. The main reason to include GH prior is that this prior is not fully included in our chosen class of priors, and hence any optimality property of it does not follow theoretically using our results. Hence, it is of interest to also study the simulation performance of their decision rule. Our comparison will be based on the average value of the proportion of misclassified hypotheses, which is used as an estimate of the risk with respect to additive 010-1 loss.
    The data generating scheme is as follows. For any fixed level of sparsity $p\in(0,1)$, we generate $n=200$ independent observations $Y_{1},\cdots,Y_{n}$, from a two-group model (4). Motivated by [Datta and Dunson (2016)], we take $\alpha=0.5,\beta=1,\delta=10$. All the rules under study are applied on the generated dataset of size $200$. This procedure is repeated $1000$ times and the average number of misclassifications is reported. For the simulation study, we use three parameter beta normal (TPBN) prior with both hyperparameters being $1.5$ and choose the global parameter $\tau$ of the same order as $p$. We also use the same GH prior as in the study of [Datta and Dunson (2016)]. We also use the empirical Bayes estimate of $\tau$, given in (17). The results are presented in Table 1, where $p\in\{0.01,0.02,\cdots,0.2\}$. In the table the column TPBN-tun refers to the results obtained for rule (16), while TPBN-EB refers to those for rule (18).
Table 1: Average misclassification probabilities based on 1000 replications
Sparsity level Two-group prior TPBN-tun TPBN-EB Gauss Hypergeometric
0.01 0.032 0.032 0.032 0.032
0.02 0.048 0.048 0.048 0.048
0.03 0.056 0.056 0.058 0.058
0.04 0.061 0.063 0.064 0.062
0.05 0.065 0.068 0.069 0.067
0.06 0.076 0.081 0.084 0.082
0.07 0.084 0.090 0.093 0.090
0.08 0.089 0.096 0.099 0.095
0.09 0.093 0.101 0.104 0.100
0.10 0.095 0.104 0.107 0.102
0.11 0.102 0.112 0.116 0.111
0.12 0.105 0.116 0.121 0.115
0.13 0.109 0.122 0.130 0.120
0.14 0.113 0.130 0.139 0.128
0.15 0.115 0.134 0.142 0.134
0.16 0.119 0.139 0.147 0.138
0.17 0.125 0.147 0.155 0.149
0.18 0.125 0.150 0.160 0.151
0.19 0.128 0.155 0.168 0.155
0.20 0.131 0.161 0.172 0.162
When the level of sparsity is small, the misclassification probabilities of decision rules based on our chosen one-group prior are quite close to those of the Oracle rule, which aligns well with the results established in this paper. The results corresponding to the GH prior are also similar to those of ours, which indicates that the theoretical optimality property is possibly also true for the testing rule of [Datta and Dunson (2016)].
6 Real data analysis
We also apply our modelling in the Global Terrorism Dataset for the years 1970-2020, which consists of the number of terrorist attacks in different countries over the previously mentioned period. This dataset is available on the website Global Terrorism Database on request. The following plot contains the number of terrorist attacks reported in different countries (names blinded) over the years, along with corresponding estimates using our approach and that of the GH prior-based approach. The estimates are based on the posterior expected means of the rates of the terrorist attacks for different countries. We have used the empirical Bayes version of the posterior mean here as the fraction of “nulls” is unknown here and used $\widehat{\tau}=\max\bigl{\{}\frac{1}{n},\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{Y_{i}\geq{100}\}\bigr{\}}$. The change in the definition of $\widehat{\tau}$ from (18) is done due to the fact that we found it is reasonable to consider an observation coming from “non-null” group if the average number of attacks in one year is at least $2$. It is quite clear from the figure that our prior provides a better estimates than those of the GH prior when the number of terrorist attacks is moderate.
Figure 1: Performance of our method on real dataset
On the other hand, even if one is interested in considering the countries which are worst hit, then also, our approach provides a more accurate estimate of the true number of attacks than that of [Datta and Dunson (2016)]. The counts for the top 10 worst hit countries along with the estimates obtained by us and that using [Datta and Dunson (2016)] are illustrated in Table 2.
Table 2: Performance of our method on real dataset for worst hit countries
Country no. No. of Terrorist Attacts Estimate using GH Estimate using TPBN
Country 1 426 424 424
Country 2 386 384 385
Country 3 334 331 332
Country 4 311 307 309
Country 5 249 244 248
Country 6 240 235 239
Country 7 216 210 215
Country 8 225 220 224
Country 9 198 191 197
Country 10 198 191 196
Finally, in order to understand the role of estimate of $\tau$ based on the dataset, we consider different empirical Bayes estimates of $\tau$. We consider countries with total number of attacks at least $100$ as belonging to non-null group. Here the difference in estimates lies in using different thresholds to categorize an observation as a signal or not. We choose three values of the thresholds, namely $50,100$ and $200$, namely, and compute the misclassification probabilities. The results are provided in Table 3. It is clear from the table that the choice of $\tau$ plays a pivotal role in controlling the misclassification probability. The misclassification probability, when the sample estimate of $\tau$ is of order $p$, is significantly less compared to other choices of $\tau$. This provides a clear picture of the role of the global parameter $\tau$, when in a multiple testing problem involving count data, the mean parameter is modeled by one-group prior.
Table 3: Misclassification probabilities with the choice of empirical Bayes estimate of $\tau$
Non-null Estimate of $\tau$ MP
$\{Y_{i}\geq 100\}$ $\max\bigl{\{}\frac{1}{n},\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{Y_{i}\geq{50}\}\bigr{\}}$ 0.09448
$\{Y_{i}\geq 100\}$ $\max\bigl{\{}\frac{1}{n},\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{Y_{i}\geq{100}\}\bigr{\}}$ 0.00787
$\{Y_{i}\geq 100\}$ $\max\bigl{\{}\frac{1}{n},\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{Y_{i}\geq{200}\}\bigr{\}}$ 0.02362
7 Concluding remarks and scope for future work
In this paper, we study modelling of high-dimensional count data containing a lot of values near zero, namely “quasi-sparse” count data. Assuming that observations are generated from a Poisson distribution with mean $\theta_{i}$ and modelling $\theta_{i}$ by (3), at first, we obtained an expression for the optimal Bayes risk under an additive 010-1 loss function. This was obtained under suitable assumptions on the model parameters. Next, our focus was to study the optimality property of a decision rule based on a class of one-group priors when the true data is generated from a two-group model. In this context, we have been able to establish that irrespective of whether the level of sparsity is known or not, our decision rules using global-local priors attain the optimal Bayes risk, upto some multiplicative constant. To the best of our knowledge, these results are the first of their kind in the literature on sparse count data.
    Throughout our calculations, we either treat $\tau$ as a tuning parameter or use an empirical Bayes estimate of it. A question that arises naturally is whether the same optimality property holds for a similar decision rule when one uses a non-degenerate prior on $\tau$. Questions can also be asked regarding the choice of the prior on $\tau$ or the range on which it has to be defined. Though we have not discussed these here, we hope techniques used in [Paul and Chakrabarti (2025)] might be helpful for this purpose also.
    [Hamura et al. (2022)] also studied a class of global-local priors for modeling quasi-sparse count data. They proposed two priors, namely, inverse gamma (IG) prior and extremely heavy-tailed (EH) prior, and proved that both of them possess the tail robustness property introduced by [Carvalho et al. (2010)]. However, they kept the global parameter fixed and did not study the asymptotic optimality in terms of the Bayes risk of the decision rule induced by their chosen class of priors. We expect that by letting the global parameter go to zero as the sample size increases and using arguments similar to ours, one may be able to establish results similar to Theorem 1 and Theorem 2 of this work. We also expect that our techniques will be handy in the case of the Gauss Hypergeometric prior used by [Datta and Dunson (2016)] for providing a non-trivial bound on the probability of type II error and the corresponding Bayes risk of their decision rule, irrespective of the underlying sparsity to be known or unknown. It may be recalled that the simulation performance of the GH prior is very encouraging in terms of Bayes risk.
    As mentioned by [Hamura et al. (2022)], one possible drawback of the Poisson modelling is that for the Poisson distribution, the mean and the variance are the same, a situation which does not always hold, specially in sparse count data. One possible remedy is to use the Negative Binomial distribution to model the situation and try to provide answer to the same questions discussed here. However, instead of considering a particular distribution, one may consider a general class of distributions and study such optimality for that class. We have already obtained some results in this context when observations are generated from a subclass of one parameter exponential family distributions containing Normal, Poisson and negative binomial as special cases. Interestingly, these three also belong to the Natural Exponential Family with Quadratic Variance function, or in short, NEF-QVF, proposed by [Morris (1982)]. However, optimality results in terms of Bayes risk are still not established for the class of [Morris (1982)]. We will be actively pursuing this in the near future.
8 Proofs
Lemma 17.
If $L$ is a slowly varying function L:
  1. 1.
    $L^{\alpha}$ is slowly varying for all $\alpha\in\mathbb{R}$.
  2. 2.
    $\frac{\log L(x)}{\log x}\to 0$ as $x\to\infty$.
  3. 3.
    for every $\alpha>0,\ x^{-\alpha}L(x)\to 0$ and $x^{\alpha}L(x)\to\infty$ as $x\to\infty$.
  4. 4.
    for $\alpha<-1,-\frac{\int_{x}^{\infty}t^{\alpha}L(t)dt}{x^{\alpha+1}L(x)}\to\frac{1}{\alpha+1}$ as $x\to\infty$.
  5. 5.
    there exists a global constant $A_{0}>0$ such that, for any $\alpha>-1,$ $\frac{\int_{A_{0}}^{x}t^{\alpha}L(t)dt}{x^{\alpha+1}L(x)}\to\frac{1}{\alpha+1}$ as $x\to\infty.$
Lemma 18.
Let $L:(0,\infty)\to(0,\infty)$ be a measurable function satisfying Assumption 2 of Section 3 and $a$ be any positive real number. Suppose $\tau\to 0$ as $n\to\infty$. Then for any $y=0,1,2,\cdots$, $\alpha>0$, and fixed $K_{1}>K_{0}>\max\{1,t_{0}\}$,
$\int_{0}^{1}u^{a+\alpha-1}(1-u)^{y-a-1}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{u}-1)\bigg{)}du\geq\frac{c_{0}(K_{0}^{-a}-K_{1}^{-a})}{a}(\tau^{2})^{y-a}(1+K_{1}\tau^{2})^{-(y+\alpha)}\hskip 1.42271pt,$
where $t_{0}$ is as in Assumption 2 of Section 3.
Proof.
Let $I=\int_{0}^{1}u^{a+\alpha-1}(1-u)^{y-a-1}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{u}-1)\bigg{)}du$. Then with the change of variable $t=\frac{1}{\tau^{2}}(\frac{1}{u}-1)$, we have
$I=(\tau^{2})^{y-a}\int_{0}^{\infty}(1+t\tau^{2})^{-(y+\alpha)}t^{y-a-1}L(t)dt\hskip 1.42271pt.$ (22)
By Assumption 2 on $L(\cdot)$, there exists $K_{0}(>1)$ such that $L(t)\geq c_{0}$ if $t\geq K_{0}\geq t_{0}$. Hence, we obtain
$\int_{0}^{\infty}(1+t\tau^{2})^{-(y+\alpha)}t^{y-a-1}L(t)dt$ $\geq\int_{K_{0}}^{\infty}(1+t\tau^{2})^{-(y+\alpha)}t^{y-a-1}L(t)dt$
$\geq c_{0}\int_{K_{0}}^{\infty}(1+t\tau^{2})^{-(y+\alpha)}t^{-a-1}dt$
$\geq c_{0}\int_{K_{0}}^{K_{1}}(1+t\tau^{2})^{-(y+\alpha)}t^{-a-1}dt$
$\geq\frac{c_{0}(K_{0}^{-a}-K_{1}^{-a})}{a}(1+K_{1}\tau^{2})^{-(y+\alpha)}.$ (23)
Finally, the result follows from (22) and (8). ∎
Proof of Theorem 3.
Fix any $\epsilon\in(0,1)$. Then
$\mathbb{P}(\kappa_{i}<\epsilon|Y_{i},\tau)$ $=\frac{\int_{0}^{\epsilon}{\kappa_{i}^{a+\alpha-1}}(1-\kappa_{i})^{Y_{i}-a-1}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)\bigg{)}d\kappa_{i}}{\int_{0}^{1}\kappa_{i}^{a+\alpha-1}(1-\kappa_{i})^{Y_{i}-a-1}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)\bigg{)}d\kappa_{i}}\hskip 1.42271pt.$
Now using the change of variable $t=\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)$ to the numerator and denominator of the previous quantity and applying (8) of Lemma 18 for the denominator, we have
$\mathbb{P}(\kappa_{i}<\epsilon|Y_{i},\tau)\leq\frac{a}{c_{0}}(K_{0}^{-a}-K_{1}^{-a})^{-1}(1+K_{1}\tau^{2})^{(Y_{i}+\alpha)}\int_{\frac{1}{\tau^{2}}(\frac{1}{\epsilon}-1)}^{\infty}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt\hskip 1.42271pt.$ (24)
Also, note that
$\int_{\frac{1}{\tau^{2}}(\frac{1}{\epsilon}-1)}^{\infty}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt$ $=(\tau^{2})^{-Y_{i}-\alpha}\int_{\frac{1}{\tau^{2}}(\frac{1}{\epsilon}-1)}^{\infty}\bigg{(}\frac{t\tau^{2}}{1+t\tau^{2}}\bigg{)}^{Y_{i}+\alpha}t^{-\alpha-a-1}L(t)dt$
$\leq(\tau^{2})^{-Y_{i}-\alpha}\int_{\frac{1}{\tau^{2}}(\frac{1}{\epsilon}-1)}^{\infty}t^{-\alpha-a-1}L(t)dt\hskip 1.42271pt.$ (25)
Next using Lemma 17 and the definition of $L(\cdot)$, we obtain
$\int_{\frac{1}{\tau^{2}}(\frac{1}{\epsilon}-1)}^{\infty}t^{-\alpha-a-1}L(t)dt=\frac{1}{(a+\alpha)}(\tau^{2})^{a+\alpha}\bigg{(}\frac{\epsilon}{1-\epsilon}\bigg{)}^{a+\alpha}L\bigg{(}\frac{1}{\tau^{2}}\bigg{)}(1+o(1))\hskip 1.42271pt.$ (26)
Using (24)-(26) the proof of Theorem 3 is obtained easily. ∎
Proof of Theorem 4.
Using the definition,
$\mathbb{E}(1-\kappa_{i}|Y_{i},\tau)=\frac{\int_{0}^{1}{\kappa_{i}^{a+\alpha-1}}(1-\kappa_{i})^{Y_{i}-a}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)\bigg{)}d\kappa_{i}}{\int_{0}^{1}\kappa_{i}^{a+\alpha-1}(1-\kappa_{i})^{Y_{i}-a-1}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)\bigg{)}d\kappa_{i}}\hskip 1.42271pt.$
Again using the change of variable $t=\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)$ in the numerator and denominator of the previous quantity and applying (8) of Lemma 18 for the denominator, we have
$\mathbb{E}(1-\kappa_{i}|Y_{i},\tau)\leq\frac{a}{c_{0}}(K_{0}^{-a}-K_{1}^{-a})^{-1}\tau^{2}(1+K_{1}\tau^{2})^{(Y_{i}+\alpha)}\int_{0}^{\infty}(1+t\tau^{2})^{-(Y_{i}+\alpha+1)}t^{Y_{i}-a}L(t)dt\hskip 1.42271pt.$ (27)
Next, we divide the range of integration into two parts, namely in $t\in(0,1)$ and $t\geq 1$. We observe that
$\tau^{2}\int_{0}^{1}(1+t\tau^{2})^{-(Y_{i}+\alpha+1)}t^{Y_{i}-a}L(t)dt\leq\tau^{2}K^{-1}\hskip 1.42271pt.$ (28)
Next using Assumption 2 on $L(\cdot)$, it is easy to see that for any $Y_{i}\in[0,a-1)$,
$\tau^{2}\int_{1}^{\infty}(1+t\tau^{2})^{-(Y_{i}+\alpha+1)}t^{Y_{i}-a}L(t)dt\leq\frac{\tau^{2}M}{(a-Y_{i}-1)}\hskip 1.42271pt.$ (29)
Finally, Theorem 4 follows from (27)-(29). ∎
Proof of Theorem 6.
Fix any $\eta\in(0,1)$ and $\delta_{1}\in(0,1)$. Now using the definition,
$\mathbb{P}(\kappa_{i}>\eta|Y_{i},\tau)=\frac{\int_{\eta}^{1}\kappa_{i}^{a+\alpha-1}(1-\kappa_{i})^{Y_{i}-a-1}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)\bigg{)}d\kappa_{i}}{\int_{0}^{1}\kappa_{i}^{a+\alpha-1}(1-\kappa_{i})^{Y_{i}-a-1}L\bigg{(}\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)\bigg{)}d\kappa_{i}}\hskip 1.42271pt.$
Then using the change of variable $t=\frac{1}{\tau^{2}}(\frac{1}{\kappa_{i}}-1)$ to both the numerator and denominator of the right-hand side of the above equation, we obtain, for any $\delta_{1}\in(0,1)$,
$\mathbb{P}(\kappa_{i}>\eta|Y_{i},\tau)=\frac{\int_{0}^{\frac{1}{\tau^{2}}(\frac{1}{\eta}-1)}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt}{\int_{0}^{\infty}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt}$
$\leq\frac{\int_{0}^{\frac{1}{\tau^{2}}(\frac{1}{\eta}-1)}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt}{\int_{\frac{1}{\tau^{2}}(\frac{1}{\eta\delta_{1}}-1)}^{\infty}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt}\hskip 1.42271pt.$ (30)
Note that the numerator of (30) can be further bounded as
$\int_{0}^{\frac{1}{\tau^{2}}(\frac{1}{\eta}-1)}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt$ $=(\tau^{2})^{-Y_{i}}\int_{0}^{\frac{1}{\tau^{2}}(\frac{1}{\eta}-1)}(\frac{t\tau^{2}}{1+t\tau^{2}})^{Y_{i}}(1+t\tau^{2})^{-\alpha}t^{-a-1}L(t)dt$
$\leq(\tau^{2})^{-Y_{i}}(1-\eta)^{Y_{i}}\int_{0}^{\infty}t^{-a-1}L(t)dt$
$=K^{-1}\bigg{(}\frac{1-\eta}{\tau^{2}}\bigg{)}^{Y_{i}}\hskip 1.42271pt.$ (31)
Here the inequality occurs due to the fact that $t\tau^{2}/(1+t\tau^{2})$ is increasing in $t$ for any fixed $\tau>0$ whenever $t\in(0,\frac{1}{\tau^{2}}(\frac{1}{\eta}-1))$. Next we use $\int_{0}^{\infty}t^{-a-1}L(t)dt=K^{-1}$.
On the other hand, for the denominator, we have
$\int_{\frac{1}{\tau^{2}}(\frac{1}{\eta\delta_{1}}-1)}^{\infty}(1+t\tau^{2})^{-(Y_{i}+\alpha)}t^{Y_{i}-a-1}L(t)dt$
$=(\tau^{2})^{-Y_{i}-\alpha}\int_{\frac{1}{\tau^{2}}(\frac{1}{\eta\delta_{1}}-1)}^{\infty}\bigg{(}\frac{t\tau^{2}}{1+t\tau^{2}}\bigg{)}^{Y_{i}+\alpha}t^{-a-\alpha-1}L(t)dt$
$\geq(\tau^{2})^{-Y_{i}-\alpha}(1-\eta\delta_{1})^{Y_{i}+\alpha}\int_{\frac{1}{\tau^{2}}(\frac{1}{\eta\delta_{1}}-1)}^{\infty}t^{-a-\alpha-1}L(t)dt$
$\geq c_{0}(\tau^{2})^{-Y_{i}-\alpha}(1-\eta\delta_{1})^{Y_{i}+\alpha}\int_{\frac{1}{\tau^{2}}(\frac{1}{\eta\delta_{1}}-1)}^{\infty}t^{-a-\alpha-1}dt$
$=\frac{c_{0}}{(a+\alpha)}\bigg{(}\frac{1}{\tau^{2}}\bigg{)}^{Y_{i}-a}(1-\eta\delta_{1})^{Y_{i}-a}(\eta\delta_{1})^{a+\alpha}\hskip 1.42271pt.$ (32)
In the chain of inequalities, the first one holds due to the fact that $t\tau^{2}/(1+t\tau^{2})$ is increasing in $t$ for any fixed $\tau>0$ whenever $t\geq\frac{1}{\tau^{2}}(\frac{1}{\eta\delta_{1}}-1)$. The second inequality follows from Assumption 2 on $L(\cdot)$ by noting that for any fixed $\eta\in(0,1)$ and $\delta_{1}\in(0,1)$, $\tau$ can be made small enough so that $\frac{1}{\tau^{2}}(\frac{1}{\eta\delta_{1}}-1)\geq t_{0}$. Finally, (8) and (8) provide the upper bound on (30) and complete the proof of Theorem 6. ∎
Proof of Theorem 10.
In order to calculate type I error, note that, the posterior distribution of $\kappa_{i}$ given $Y_{i}$ and $\tau$ has the same from for all $i=1,2,\cdots,n$. Also, under $H_{0i}$, the marginal distribution of $Y_{i}$ is the same for all $i=1,2,\cdots,n$. These two facts indicate that the probability of type I error of $i^{\text{th}}$ decision rule, denoted as $t_{1i}$ is independent of $i$. We further then have,
$t_{1i}\equiv t_{1}$ $=\mathbb{P}_{H_{0i}}({E}(1-\kappa_{i}|Y_{i},\tau)>\frac{1}{2})$
$\leq\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\tau)>\frac{1}{2},Y_{i}\leq\frac{a}{2})+\mathbb{P}_{H_{0i}}(Y_{i}>\frac{a}{2})\hskip 1.42271pt.$ (33)
Now concentrate on the first term in (8). In this case, using Theorem 4, as $\tau\to 0$, this probability can be further bounded as
$\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\tau)>\frac{1}{2},Y_{i}\leq\frac{a}{2})$
$\leq\mathbb{P}_{H_{0i}}\bigg{(}M_{1}\tau^{2}[K^{-1}+\frac{M}{\frac{a}{2}-1}](1+K_{1}\tau^{2})^{(Y_{i}+\alpha)}>\frac{1}{2},Y_{i}\leq\frac{a}{2}\bigg{)}$
$\leq\mathbb{P}\bigg{(}M_{2}\tau^{2}(1+K_{1}\tau^{2})^{(\frac{a}{2}+\alpha)}>\frac{1}{2}\bigg{)}$
$=0,\textrm{for all sufficiently large }n,$ (34)
where $M_{1}$ and $M_{2}$ are constants depending only on $c_{0},a$, $\alpha,K_{0}$ and $K_{1}$ and is independent of $Y_{i}$. Since, under $H_{0i}$, $Y_{i}\stackrel{{\scriptstyle iid}}{{\sim}}NB(\alpha,\frac{1}{\beta+1})$, for the second term of (8), using the Markov’s inequality, we have
$\mathbb{P}_{H_{0i}}(Y_{i}>\frac{a}{2})\leq\frac{2\alpha\beta}{a}\hskip 1.42271pt.$ (35)
The proof is completed using (8)-(35). ∎
Proof of Theorem 12.
Using exactly a similar argument as used in Theorem 6 about the equality of $t_{1i}$ for $i=1,2,\cdots,n$, in this case also, under $H_{1i}$, the distribution of $Y_{i}$ is the same for all $i$. Also, note that, the form of the posterior distribution of $\kappa_{i}$ given $Y_{i}$ and $\tau$ is same for all $i=1,2,\cdots,n$. As a consequence of this, the probability of type II error of $i^{\text{th}}$ decision rule is denoted as $t_{2}$.
Note that, using Theorem 6, for any given $\eta\in(0,1)$ and $\tau>0$, $E(\kappa_{i}|Y_{i},\tau)$ can be bounded as,
$\mathbb{E}(\kappa_{i}|Y_{i},\tau)$ $=E(\kappa_{i}1\{\kappa_{i}\leq\eta\}|Y_{i},\tau)+E(\kappa_{i}1\{\kappa_{i}>\eta\}|Y_{i},\tau)$
$\leq\eta+\mathbb{P}(\kappa_{i}>\eta|Y_{i},\tau)$
$\leq\eta+\frac{1}{\tilde{K}}\tau^{-2a}r^{Y_{i}},$ (36)
where $\tilde{K}$ is a constant depending on $c_{0},K,\eta$ and $\delta_{1}$ and $r=\frac{1-\eta}{1-\eta\delta_{1}}$. Hence, for any $\eta\in(0,\frac{1}{2})$, for all sufficiently large $n$, the upper bound on $t_{2}$ is of the form
$t_{2}$ $=\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},\tau)\geq\frac{1}{2})$
$\leq\mathbb{P}_{H_{1i}}\bigg{(}\tau^{-2a}r^{Y_{i}}\geq\tilde{K}_{1}\bigg{(}\frac{1}{2}-\eta\bigg{)}\bigg{)}$
$=\mathbb{P}_{H_{1i}}(Y_{i}\log r-2a\log\tau\geq K_{2})$
$=\mathbb{P}_{H_{1i}}\bigg{(}-Y_{i}\log(\frac{1}{r})\geq K_{2}-2a\log(\frac{1}{\tau})\bigg{)}$
$=\mathbb{P}_{H_{1i}}\bigg{(}Y_{i}\leq\frac{2a}{\log(\frac{1}{r})}\log(\frac{1}{\tau})(1+o(1))\bigg{)}$
$=\mathbb{P}_{H_{1i}}\bigg{(}Y_{i}\leq\frac{2a}{\log(\frac{1}{r})}\log(\frac{1}{p})(1+o(1))\bigg{)}$
$=\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq(\frac{2a}{\sqrt{\alpha}\log(\frac{1}{r})}\cdot\frac{\log(\frac{1}{p})}{\delta}-\sqrt{\alpha})(1+o_{2}(1))\bigg{)}$
$=\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq-\sqrt{\alpha})(1+o_{2}(1))\bigg{)}\hskip 1.42271pt,$ (37)
where the $o(1),o_{2}(1)$ terms are non-random, independent of any $i$ and tend to zero as $n\to\infty$. In the chain of equalities, the penultimate step follows using the assumption that $\lim_{n\to\infty}\tau/p\in(0,\infty)$. Here $Z_{i}$ is the standardized version of $Y_{i}$ under $H_{1i}$. This completes the proof of Theorem 12.
Proof of Theorem 14.
First, let us find an expression for $\gamma_{n}=P(Y_{i}\geq 1)$. We have,
$\gamma_{n}$ $=(1-p)\mathbb{P}_{H_{0i}}(Y_{i}\geq 1)+p\mathbb{P}_{H_{1i}}(Y_{i}\geq 1)$
$=p[\mathbb{P}_{H_{1i}}(Y_{i}\geq 1)+\frac{(1-p)}{p}\mathbb{P}_{H_{0i}}(Y_{i}\geq 1)]$
Since, $Y_{i}\stackrel{{\scriptstyle iid}}{{\sim}}NB(\alpha,\frac{1}{\beta+1})$, under $H_{0i}$, using Markov’s inequality and Assumption 1, we have, $\frac{1}{p}\mathbb{P}_{H_{0i}}(Y_{i}\geq 1)\to 0$ as $n\to\infty$. Also, note that, under $H_{1i}$,
$\mathbb{P}(Y_{i}\geq 1)=(1-(\beta+\delta+1)^{-\alpha}),$
which tends to $1$ under Assumption 1. Combining these two, we obtain, as $n\to\infty$, $\gamma_{n}=(1-(\beta+\delta+1)^{-\alpha})p(1+o(1))$.
Now, using the definition of type I error of the $i^{\text{th}}$ decision rule in (18), we have
$t^{\text{EB}}_{1i}$ $=\mathbb{P}_{H_{0i}}(\mathbb{E}(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2})$
$=\mathbb{P}_{H_{0i}}(\mathbb{E}(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},0\leq Y_{i}\leq\frac{a}{2})+\mathbb{P}_{H_{0i}}(\mathbb{E}(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},Y_{i}>\frac{a}{2})\hskip 1.42271pt.$ (38)
For the second term of (8), using Markov’s inequality, we have
$\mathbb{P}_{H_{0i}}(\mathbb{E}(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},Y_{i}>\frac{a}{2})\leq\mathbb{P}_{H_{0i}}(Y_{i}>\frac{a}{2})\leq\frac{2\alpha\beta}{a}\hskip 1.42271pt.$ (39)
For the first term in (8), we obtain
$\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},0\leq Y_{i}\leq\frac{a}{2})$
$=\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},0\leq Y_{i}\leq\frac{a}{2},\widehat{\tau}\leq 2\gamma_{n})$
$+\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},0\leq Y_{i}\leq\frac{a}{2},\widehat{\tau}>2\gamma_{n})\hskip 1.42271pt.$ (40)
Now, using the form of $\gamma_{n}$, we note that
$\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},0\leq Y_{i}\leq\frac{a}{2},\widehat{\tau}\leq 2\gamma_{n})$
$\leq\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},2\gamma_{n})>\frac{1}{2},0\leq Y_{i}\leq\frac{a}{2})$
$=0,\textrm{ for all sufficiently large }n.$ (41)
The inequality above holds due to the fact that for any fixed $y\geq 0$, $E(1-\kappa|y,\tau)$ is non-decreasing in $\tau$. The last assertion follows from (8).
  Now let us concentrate on the second term of (8). Define $\widehat{\tau}_{1}=\frac{1}{n}$ and $\widehat{\tau}_{2}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{Y_{i}\geq 1\}$. Hence, $\widehat{\tau}=\text{max}\{\widehat{\tau}_{1},\widehat{\tau}_{2}\}$. Also, note that $\gamma_{n}\sim(1-(\beta+\delta+1)^{-\alpha})p$ and $p\propto n^{-\epsilon},0<\epsilon<1$ implies $\frac{1}{n}<2\gamma_{n}$ for all sufficiently large $n$. Next, we use the fact that $\{\widehat{\tau}>2\gamma_{n}\}$ implies that either of the two events $\{\widehat{\tau}_{1}>2\gamma_{n}\}$ or $\{\widehat{\tau}_{2}>2\gamma_{n}\}$ will happen. Using these facts for all sufficiently large $n$, we have
$\mathbb{P}_{H_{0i}}(E(1-\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},0\leq Y_{i}\leq\frac{a}{2},\widehat{\tau}>2\gamma_{n})\leq\mathbb{P}_{H_{0i}}(\widehat{\tau}>2\gamma_{n})$
$\leq\mathbb{P}_{H_{0i}}(\widehat{\tau}_{1}>2\gamma_{n})+\mathbb{P}_{H_{0i}}(\widehat{\tau}_{2}>2\gamma_{n})$
$=\mathbb{P}_{H_{0i}}(\widehat{\tau}_{2}>2\gamma_{n})$
$\leq\mathbb{P}_{H_{0i}}(\widehat{\tau}_{2}>2\gamma_{n},Y_{i}<1)+\mathbb{P}_{H_{0i}}(Y_{i}\geq 1)$
$\leq\mathbb{P}_{H_{0i}}(\widehat{\tau}_{2}>2\gamma_{n},Y_{i}<1)+{\alpha\beta}{}\hskip 1.42271pt.$ (42)
Now we are only left with only the first term of (8). Note that
$\{\widehat{\tau}_{2}>2\gamma_{n},Y_{i}<1\}\subseteq\{\frac{1}{n}\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}\mathbf{1}\{Y_{j}\geq 1\}>2\gamma_{n}\}.$
As a result, using the Chernoff-Hoeffding inequality, we have
$\mathbb{P}_{H_{0i}}(\widehat{\tau}_{2}>2\gamma_{n},Y_{i}<1)$ $\leq\mathbb{P}_{H_{0i}}(\frac{1}{n}\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}\mathbf{1}\{Y_{j}\geq 1\}>2\gamma_{n})$
$\leq\mathbb{P}(\frac{1}{n-1}\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}\mathbf{1}\{Y_{j}\geq 1\}\geq 2\gamma_{n})$
$\leq e^{-(n-1)D(2\gamma_{n}||\gamma_{n})}\hskip 1.42271pt,$ (43)
where $D(x||y)$ is defined as, $D(x||y)=x\log(\frac{x}{y})+(1-x)\log(\frac{1-x}{1-y})$. Next using these facts that $\log(\frac{1}{1-x})\sim x$ as $x\to 0$ and $\gamma_{n}\to 0$ as $n\to\infty$ and also using the calculations of Theorem 10 of [Ghosh et al. (2016)], we have
$D(2\gamma_{n}||\gamma_{n})=(2\log 2-1)\gamma_{n}(1+o(1))=(2\log 2-1)(1-(\beta+\delta+1)^{-\alpha})p(1+o(1))\hskip 1.42271pt.$ (44)
Here $o(1)$ is non-random and tends to 0 as $n\to\infty$. Hence using (8) and (44), we obtain
$\mathbb{P}_{H_{0i}}(\widehat{\tau}_{2}>2\gamma_{n},Y_{i}<1)\leq e^{-(2\log 2-1)(1-(\beta+\delta+1)^{-\alpha})np(1+o(1))}\hskip 1.42271pt.$ (45)
Combining the above arguments the proof of Theorem 14 follows. ∎
Proof of Theorem 15.
Let us fix any $C_{3}\in(0,1)$. Now, using the definition of type II error of the $i^{\text{th}}$ decision rule in (18), we have
$t^{\text{EB}}_{2i}$ $=\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},\widehat{\tau})\geq\frac{1}{2})$
$=\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},\widehat{\tau})\geq\frac{1}{2},\widehat{\tau}<C_{3}\gamma_{n})+\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},\widehat{\tau})\geq\frac{1}{2},\widehat{\tau}\geq C_{3}\gamma_{n})\hskip 1.42271pt,$ (46)
where $\gamma_{n}$ is same as defined in the previous theorem. Let us first bound the second term in the r.h.s. of (8). Using the fact that for any fixed $y$, $E(\kappa|y,\tau)$ is decreasing in $\tau$, towards that, we first note that
$\{E(\kappa_{i}|Y_{i},\widehat{\tau})\geq\frac{1}{2},\widehat{\tau}\geq C_{3}\gamma_{n}\}\subseteq\{E(\kappa_{i}|Y_{i},C_{3}\gamma_{n})\geq\frac{1}{2}\}\hskip 1.42271pt.$ (47)
Now applying the same set of arguments used in the proof of Theorem 12 and noting under $H_{1i}$}, $Y_{i}\stackrel{{\scriptstyle iid}}{{\sim}}NB(\alpha,\frac{1}{\beta+\delta+1})$, we obtain, as $n\to\infty$,
$\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},C_{3}\gamma_{n})\geq\frac{1}{2})$ $\leq\mathbb{P}_{H_{1i}}\bigg{(}Y_{i}\leq\frac{2a}{\log(\frac{1}{r})}\log(\frac{1}{C_{3}\gamma_{n}})(1+o(1))\bigg{)}$
$=\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq(\frac{2a}{\sqrt{\alpha}\log(\frac{1}{r})}\cdot\frac{\log(\frac{1}{p})}{\delta}-\sqrt{\alpha})(1+o_{2}(1))\bigg{)}$
$=\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq-\sqrt{\alpha})(1+o_{2}(1))\bigg{)}\hskip 1.42271pt,$ (48)
where $Z_{i}$ is the standardized version of $Y_{i}$ under $H_{1i}$ and the term $o_{2}(1)$ is non-random, independent of any $i$ and tends to zero as $n\to\infty$. Here we first use $\gamma_{n}=(1-(\beta+\delta+1)^{-\alpha})p(1+o(1))$ and next use the same argument used to find an upper bound of $t_{2}$ in Theorem 12.
Now our aim is to show that the first term in (8) goes to 0 as $n\to\infty$. Note that, $\widehat{\tau}_{2}\geq\frac{1}{n}\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}\mathbf{1}\{Y_{j}\geq 1\}$. Using this observation and noting that due to independence, the distribution of remaining $Y_{j}$’s do not depend on that of $Y_{i}$, we have
$\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},\widehat{\tau})\geq\frac{1}{2},\widehat{\tau}<C_{3}\gamma_{n})$ $\leq\mathbb{P}_{H_{1i}}(\widehat{\tau}<C_{3}\gamma_{n})$
$\leq\mathbb{P}_{H_{1i}}(\widehat{\tau}_{2}<C_{3}\gamma_{n})$
$\leq\mathbb{P}_{H_{1i}}(\frac{1}{n}\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}\mathbf{1}\{Y_{j}\geq 1\}\leq C_{3}\gamma_{n})$
$=\mathbb{P}(-\frac{1}{n-1}(\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}(\mathbf{1}\{Y_{j}\geq 1\}-\gamma_{n}))\geq(1-\frac{n}{n-1}C_{3})\gamma_{n})\hskip 1.42271pt,$ (49)
where probability in the last statement is calculated using the marginal distribution (4) of $Y_{j},j\neq i$. Since, $1-\frac{n}{n-1}C_{3}\to 1-C_{3}$ as $n\to\infty$ and $C_{3}\in(0,1)$, so, $1-\frac{n}{n-1}C_{3}>0$ for all sufficiently large $n$. Finally using Markov’s inequality,
$\mathbb{P}(-\frac{1}{n-1}(\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}(\mathbf{1}\{Y_{i}\geq 1\}-\gamma_{n}))\geq(1-\frac{n}{n-1}C_{3})\gamma_{n})$
$\leq\mathbb{P}\bigg{(}|\frac{1}{n-1}(\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{n}(\mathbf{1}\{Y_{i}\geq 1\}-\gamma_{n}))|\geq(1-\frac{n}{n-1}C_{3})\gamma_{n}\bigg{)}$
$\leq\frac{\gamma_{n}(1-\gamma_{n})}{(n-1)(1-\frac{n}{n-1}C_{3})^{2}\gamma^{2}_{n}}$
$=\frac{(1-\gamma_{n})}{(1-C_{3})^{2}n\gamma_{n}}(1+o(1))=o(1)\hskip 1.42271pt.$ (50)
At the final step, we use $\gamma_{n}\sim(1-(\beta+\delta+1)^{-\alpha})p$ and $p\propto n^{-\epsilon},0<\epsilon<1$. Combining (8) and (8), we obtain, as $n\to\infty$,
$\mathbb{P}_{H_{1i}}(E(\kappa_{i}|Y_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}<C_{3}\gamma_{n})=o(1)\hskip 1.42271pt.$
Hence using (8), as $n\to\infty$, we have
$t^{\text{EB}}_{2i}\leq\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq-\sqrt{\alpha})(1+o_{2}(1))\bigg{)}(1+o(1))\hskip 1.42271pt,$ (51)
since the r.h.s. of (8) is bounded away from zero under Assumption 1. This completes the proof of Theorem 15. ∎
Proof of Theorem 1.
First, we study the probabilities of type I error ($t^{\text{BO}}_{1}$) and type II error ($t^{\text{BO}}_{1}$) of the decision rule when each $\theta_{i}$ is modeled by a two-group prior of the form (3).
Since, under $H_{0i}$, $Y_{i}\stackrel{{\scriptstyle iid}}{{\sim}}NB(\alpha,\frac{1}{\beta+1})$, using Markov’s inequality and Assumption 1, $t^{\text{BO}}_{1}$ can be bounded as
$t^{\text{BO}}_{1}$ $=\mathbb{P}_{H_{0i}}(Y_{i}>C_{p,\alpha,\beta,\delta})$
$=\mathbb{P}_{H_{0i}}\big{(}Y_{i}>\frac{\log(\frac{1}{p})(1+\alpha\tilde{C})}{\log(\frac{1}{\beta})}(1+o(1))\big{)}$
$\leq\alpha\cdot\frac{\beta\log(\frac{1}{\beta})}{\log(\frac{1}{p})}(1+o(1))\hskip 1.42271pt.$ (52)
Since under the assumption $\beta\propto p^{C_{1}}$, for some $C_{1}>1$, we have, $\frac{t^{\text{BO}}_{1}}{p}\to 0$ as $n\to\infty$. Again using Assumption 1, we have
$t^{\text{BO}}_{2}$ $=\mathbb{P}_{H_{1i}}(Y_{i}\leq C_{p,\alpha,\beta,\delta})$
$=\mathbb{P}_{H_{1i}}\bigg{(}Z_{i}\leq\frac{C_{p,\alpha,\beta,\delta}}{\delta}-\sqrt{\alpha}(1+o_{1}(1))\bigg{)}$
$=\mathbb{P}_{H_{1i}}(Z_{i}\leq-\sqrt{\alpha}(1+o_{1}(1)))\hskip 1.42271pt.$ (53)
Note that, the equality in the second line is obtained by subtracting the mean of $Y_{i}$ and dividing by its standard deviation under $H_{1i}$}. Equality in the last line holds since under Assumption 1, $\frac{C_{p,\alpha,\beta,\delta}}{\delta}\to 0$ as $n\to\infty$. Now, observe that, under Assumption 1, for any $0<\epsilon<1$, for all sufficiently large $n$, the power of the Bayes Oracle rule, denoted as ${Power}^{\text{BO}}_{\text{Opt}}$, can be bounded as follows:
${Power}^{\text{BO}}_{\text{Opt}}$ $=\mathbb{P}_{H_{1i}}(Z_{i}>-\sqrt{\alpha}(1+o_{1}(1)))$
$\leq\mathbb{P}_{H_{1i}}(|Z_{i}|>|-\sqrt{\alpha}+o_{1}(1)|)$
$\leq\mathbb{P}_{H_{1i}}(|Z_{i}|>|(-1+\epsilon)\sqrt{\alpha}|)$
$\leq\frac{1}{\alpha(-1+\epsilon)^{2}}.$
Here in the chain of inequalities, the second one holds using the fact that $o_{1}(1)$ can be bounded by $\sqrt{\alpha}\epsilon$ (in absolute value), if $n$ is sufficiently large. Hence, we have, for all sufficiently large $n$,
$t^{\text{BO}}_{2}$ $\geq 1-\frac{1}{\alpha(-1+\epsilon)^{2}}.$ (54)
By choosing $\epsilon$ sufficiently small, $t^{\text{BO}}_{2}$ is bounded away from zero for all sufficiently large $n$, since $\alpha>1$. As a consequence of this, the asymptotic expression for the optimal Bayes risk based on the decision rule (8) with (9), denoted $R^{\text{BO}}_{\text{Opt}}$, can be written as
$R^{\text{BO}}_{\text{Opt}}$ $=np[\frac{(1-p)}{p}t^{\text{BO}}_{1}+t^{\text{BO}}_{2}]$
$=npt^{\text{BO}}_{2}(1+o_{1}(1))\hskip 1.42271pt.$ (55)
Now note that under Assumption 1, $t_{1}/p\to 0$ as $n\to\infty$ where $t_{1}$ is the probability of type I error corresponding to $i^{\text{th}}$ decision rule of the form (16) based on our class of one-group priors, as obtained in Theorem 10. Using this and Theorem 12, the asymptotic expression for the Bayes risk based on the decision rule, denoted by $R_{\text{OG}}$ can be written as
$R_{\text{OG}}=npt_{2}(1+o_{2}(1))\hskip 1.42271pt,$ (56)
by observing that $t_{2}$ is also bounded away from zero as in (54) using the same argument. Next, we take the ratio of (56) to (8). The lower bound is obtained using the fact that $\frac{R_{\text{OG}}}{R^{\text{BO}}_{\text{Opt}}}\geq 1$. For the upper bound, observe that, the lower bound on $t^{\text{BO}}_{2}$ is bounded away from zero and given any constant $\tilde{K}\in(\frac{1}{\alpha},1)$ with $\alpha>1$, for all sufficiently large $n$,
$\frac{R_{\text{OG}}}{R^{\text{BO}}_{\text{Opt}}}$ $\leq\frac{\tilde{K}\alpha}{\tilde{K}\alpha-1}.$
Proof of Theorem 2.
The asymptotic expression for the Bayes risk based on the decision rule (18), denoted by $R^{\text{EB}}_{\text{OG}}$, is of the form
$R^{\text{EB}}_{\text{OG}}=\sum_{i=1}^{n}[(1-p)t^{\text{EB}}_{1i}+pt^{\text{EB}}_{2i}]=p\sum_{i=1}^{n}\bigg{[}\frac{1-p}{p}t^{\text{EB}}_{1i}+t^{\text{EB}}_{2i}\bigg{]}\hskip 1.42271pt.$ (57)
Note that, we already proved in Theorem 1 that $t^{\text{BO}}_{2}=P_{H_{1i}}(Z_{i}\leq-\sqrt{\alpha}(1+o(1)))$ is bounded away from 0, for all sufficiently large $n$. Hence, using Theorem 15 along with Assumption 1, the proof of our desired result is obtained if we establish that
$\sum_{i=1}^{n}\frac{1-p}{p}t^{\text{EB}}_{1i}=o(n)\hskip 1.42271pt\text{as}\hskip 1.42271ptn\to\infty\hskip 1.42271pt.$ (58)
To prove (58), we use Theorem 14 where, it is derived that the upper bound of $t^{\text{EB}}_{1i}$ is the same for each $i=1,2,\cdots,n$. Further using the fact that $1-p\leq 1$, we obtain,
$\frac{1}{n}\sum_{i=1}^{n}\frac{1-p}{p}t^{\text{EB}}_{1i}\leq\frac{2\alpha\beta}{ap}+\frac{\alpha\beta}{p}+\frac{1}{p}e^{-(2\log 2-1)(1-(\beta+\delta+1)^{-\alpha})np(1+o(1))}\hskip 1.42271pt.$ (59)
Byg Assumption 1, the first two terms in the right-hand side of (59) go to 0 as $n\to\infty$. For the third term, note that, for $p\propto n^{-\epsilon},0<\epsilon<1$, $np\to\infty$ as $n\to\infty$ and $\log(\frac{1}{p})=o(np)$. As a result, the third term too goes to 0 as $n\to\infty$ and this proves (58). This completes the proof of Theorem 2. ∎